### v1

# Decision-dependent Robust Charging Infrastructure Planning

[…]
Decision-dependent Robust Charging Infrastructure Planning
[…]
Decision-dependent Robust Charging Infrastructure Planning for Light-duty Truck Electrification at Industrial Sites: Scheduling and Abandonment
[…]
Yifu Ding, Ruicheng Ao, Pablo Duenas-Martinez, Thomas Magnanti
[…]
Many industrial sites rely on diesel-powered light-duty trucks to transport workers and small-scale facilities, which has resulted in a significant amount of greenhouse emissions (GHGs). To address this, we developed a two-stage robust charging infrastructure planning model for electrifying light-duty trucks at industrial sites. The model is formulated as a mixed-integer linear programming (MILP) that optimizes the charging infrastructure, selected from multiple charger types and potential locations, and determines opportunity charging schedules for each truck based on the chosen infrastructure. Given the strict stopping points and schedules at industrial sites, we introduced a scheduling-with-abandonment problem, where trucks forgo charging if their waiting times exceed a maximum threshold. We also further incorporated the impacts of overnight charging and range anxiety on drivers’ waiting and abandonment behaviors. To represent stochastic, heterogeneous parking durations of trucks, we constructed a decision-dependent robust uncertainty set in which parking time variability flexibly depends on drivers’ charging choices. We applied the model in a case study of an open-pit mining site, which plans charger installations in eight zones and schedules a fleet of around 200 trucks. By decomposing the problem into monthly subproblems and using heuristic approaches, for the whole-year dataset, the model achieves an optimality gap of less than 0.1% within a reasonable computation time under diverse uncertainty scenarios.
1 Introduction
1.1 Background
Diesel-powered light-duty trucks are used for transporting and dispatching goods in urban environments and industrial sites. However, their usage could produce a significant amount of GHG emissions. In 2022, light-duty trucks were the most significant GHG emission source of transportation, contributing to 58% of the total emission of the transportation sector in the U.S. \@BBOPcite\@BAP\@BBNU.S. EPA (2024)\@BBCP. Battery electric vehicles (BEVs) play a central role in decarbonizing the transportation sector. Its associated charging infrastructure is currently categorized into two types. The first type is the level 1 and level 2 AC charging (hereafter referred to as ‘slow charging’), which generally requires between four hours and a full day to charge a vehicle battery from empty to 80% state of charge (SoC). The second type is the DC fast charging (hereafter referred to as ‘fast charging’), available at dedicated stations, which requires a relatively complicated installation process. The fast chargers can recharge an empty vehicle battery to 80% SoC in an hour \@BBOPcite\@BAP\@BBNUS Department of Transportation (2025)\@BBCP. Fast charging infrastructure also entails considerable capital investments and installation expenditures, primarily due to the power electronics required. Furthermore, its operational constraints differ from those of slow charging, as charging beyond 80% SoC is strongly discouraged for safety considerations \@BBOPcite\@BAP\@BBNMussa et al. (2017)\@BBCP. Another recharging option is the battery swapping station. The battery swapping station enables EVs to replace depleted batteries with fully charged ones swiftly \@BBOPcite\@BAP\@BBNMak et al. (2013)\@BBCP. However, due to the safety risks associated with battery swapping operations, the feasibility of deploying battery swapping stations at industrial sites should be assessed on a case-by-case basis.
In this paper, we examine the electrification of light-duty trucks at industrial sites for decarbonization. The original diesel-powered trucks are primarily employed for transporting workers and small-scale facilities. The strict operational schedules at industrial sites leave little room for trucks to change any designated parking locations or timetables. As a result, the scheduled trucks are expected to abandon charging if the anticipated waiting time for getting a charger surpasses a specified threshold. This is formulated as a ‘scheduling-with-abandonment’ problem. We conducted a case study of the open-pit mining site with 200 diesel-powered operational trucks. Without modifying any of their GPS data, we extracted their parking and driving patterns based on potential charging points and estimated their energy consumption at half-hour resolutions. Then, we developed a two-stage charging infrastructure planning and scheduling model for the entire truck fleet. The planned EV charging infrastructure satisfies the charging needs of each truck. That is meeting the energy consumption of each truck at any time, given the limited parking durations and battery energy capacity.
1.2 Related Work
1.2.1 Mixed-type charging infrastructure planning
Existing studies on charging infrastructure planning primarily focus on determining the location, number, and type of EV chargers, as well as the charging schedule using the selected installation. For electric buses and trucks at industrial sites, the charging process primarily involves opportunity or overnight charging, meaning they are only charged during their layovers and overnight parking \@BBOPcite\@BAP\@BBNAbdelwahed et al. (2020)\@BBCP. The vehicle routing is typically not considered in the electric bus scheduling, as buses usually have fixed stop points and timetables. After determining a finite set of candidate locations for potential EV charger deployment based on stop frequency, the number and types of chargers can be optimized according to trip patterns. \@BBOPcite\@BAP\@BBNHu et al. (2022)\@BBCP models a joint optimization problem for locating fast chargers at selected bus stops and optimizing schedules for opportunity charging. The model assumes that buses will be charged during layovers, and the cost takes into account the time-varying electricity price and penalty for additional waiting times. \@BBOPcite\@BAP\@BBNLegault et al. (2025)\@BBCP formulates a multi-period, integrated planning and scheduling problem. The paper models the electric bus timetable as a graph representing their charging and driving pattern, which is considered a deterministic, minimum-cost problem. \@BBOPcite\@BAP\@BBNGkiotsalitis et al. (2025)\@BBCP optimize the EV charger location and number by allocating bus trips to their nearest chargers with available time slots, assuming that buses will be driven to the optimal charger location after the trip ends.
1.2.2 EV charging scheduling with waiting time and abandonment
The scheduling problem is a fundamental and widely studied combinatorial optimization problem, with extensive applications in areas such as supply chain management, service management, and healthcare \@BBOPcite\@BAP\@BBNFonseca et al. (2024), Delarue et al. (2025)\@BBCP. EV charging can be viewed as a service that relies on a reusable resource—namely, the charging infrastructure. The EV charging scheduling problem could thus consider different attributes, such as charger availability, charging time, and waiting times for charging services. For instance, the delay or waiting time can be modeled as a monetary penalty in the objective function \@BBOPcite\@BAP\@BBNHu et al. (2022), Tao et al. (2024), Gkiotsalitis et al. (2025)\@BBCP. When the waiting time exceeds a threshold, the vehicle that committed a time-critical task or trip might abandon charging and leave to find another charging spot, which can be represented as a scheduling problem with abandonment \@BBOPcite\@BAP\@BBNG. Chen et al. (2023)\@BBCP. Another critical aspect of EV charging is range anxiety, referring to drivers’ concern about depleting their battery before reaching a charging station or their destination. Empirical studies suggest that range anxiety may incentivize the installation of additional charging stations or the adoption of larger battery capacities \@BBOPcite\@BAP\@BBNLim et al. (2015)\@BBCP. \@BBOPcite\@BAP\@BBNLu et al. (2025)\@BBCP model the impact of range anxiety by setting a threshold SoC before reaching the destination. \@BBOPcite\@BAP\@BBNTao et al. (2024)\@BBCP formulate the accumulated range anxiety level as a function of SoC depletion along the route and integrates the range anxiety into its objective function.
1.2.3 Decision-dependent robust optimization
In integrated charging infrastructure planning and scheduling problems, several uncertainties arise, including EV arrival and departure times, charging durations, and energy consumption. They can be modeled using stochastic optimization (SO), robust optimization (RO), and distributionally robust optimization (DRO). Among these formulations, RO typically ensures the highest level of reliability, making it particularly suitable for critical infrastructure planning. Different shapes of uncertainty sets, such as box or ellipse, lead to the robust counterparts and tractability \@BBOPcite\@BAP\@BBNBertsimas & Hertog (2022)\@BBCP. \@BBOPcite\@BAP\@BBNHu et al. (2022)\@BBCP model the uncertainty of passengers’ boarding and alighting time during bus trips using box and budget uncertainty sets based on the RO formulation. \@BBOPcite\@BAP\@BBNL. Chen et al. (2024)\@BBCP conduct EV charging scheduling considering the stochastic arrival time of different user groups, where an exponential cone programming formulation is introduced to approximate the upper bound of the stochastic problem. \@BBOPcite\@BAP\@BBNY. Chen & Liu (2023)\@BBCP developed a two-stage stochastic model for planning and operating shared autonomous EV systems, where scenarios represent uncertain user trip requests. An accelerated two-phase Benders decomposition-based algorithm is proposed for solving this two-stage programming problem.
However, these constructed uncertainty sets model uncertainties independent of exogenous factors such as charging decisions. A number of charging installation trials have already shown that charging decisions are very likely to impact parking and waiting durations. For instance, parking duration is strongly correlated with the charging types (e.g., whether it is slow or fast charging) \@BBOPcite\@BAP\@BBNCui et al. (2023), Zhang et al. (2025)\@BBCP. This could be modeled as the decision-dependent uncertainty set, and its uncertainty is endogenous \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP. \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP demonstrates that, in general, RO problems with decision-dependent uncertainty are NP-complete and require either global optimization or reformulation techniques such as the big-M method. Similarly, \@BBOPcite\@BAP\@BBNNoyan et al. (2022)\@BBCP introduces a decision-dependent ambiguity set in a machine scheduling problem and examines how different types of uncertainty dependence affect the model formulation. Their findings show that incorporating affine dependency on decision variables can reduce nonlinear constraints through the use of McCormick envelopes and the big-M approach.
Decision Variables User Behavior Uncertainty formulation
Location Number Scheduling Type Waiting time Range Anxiety
\@BBOPcite\@BAP\@BBNAbdelwahed et al. (2020)\@BBCP D
\@BBOPcite\@BAP\@BBNHu et al. (2022)\@BBCP RO
\@BBOPcite\@BAP\@BBNY. Chen & Liu (2023)\@BBCP SO
\@BBOPcite\@BAP\@BBNL. Chen et al. (2024)\@BBCP SO
\@BBOPcite\@BAP\@BBNTao et al. (2024)\@BBCP D
\@BBOPcite\@BAP\@BBNLu et al. (2025)\@BBCP SO
\@BBOPcite\@BAP\@BBNLegault et al. (2025)\@BBCP D
\@BBOPcite\@BAP\@BBNGkiotsalitis et al. (2025)\@BBCP D
Our paper RO
  • Notes: D = Deterministic model; RO = Robust Optimization; SO = Stochastic Optimization.
All the aforementioned work about charging infrastructure planning and scheduling is presented in Table 1.2.3. Compared to these works, our paper makes the following contributions:
  • We developed a two-stage EV charger planning model that optimizes the charger location, quantity, type, and scheduling. The model introduces a scheduling-with-abandonment formulation that reflects the industrial site’s strict operational rules and captures the impacts of overnight parking, range anxiety, and waiting times on the charging scheduling.
  • We developed a decision-dependent uncertainty set to capture the variability in parking durations and their differences across charger types (i.e., slow and fast chargers). Although decision-dependent robust optimization models are typically nonlinear, we demonstrated that the proposed affine dependence on the binary charging decision can be linearized, allowing the problem to be formulated as a MILP.
  • Considering the high computation cost of the scheduling problem over a long optimization horizon, we designed a ‘fix-and-optimize’ heuristic method to forecast parking intervals with high charging demands in the scheduling. Using the representative days or iterative planning by month, the model achieves an optimality gap of less than 0.1% within a reasonable computation time under four uncertainty scenarios.
2 Problem formulation
2.1 Problem description
The problem is formulated as a two-stage model: the first stage determines charger installations, and the second stage schedules truck charging based on the installed infrastructure. We consider a fleet of electric light-duty trucks $i\in\mathcal{I}$ operating around an industrial site based on the fixed stop points and schedules. As shown in Figure 1, a truck $i$ can be charged during its available half-hour parking time slots $\mathcal{S}_{i,d,t}$ at time $t$ on day $d$ at a charging zone $z$. One truck cannot appear in more than one zone at a time.
Figure 1: The studied multi-zone charging scheduling problem for the light-duty truck fleet; At any half-hour parking time slot $\mathcal{S}$ shown in light green, a truck can either choose to wait, abandon charging, or charge. The effective parking durations pppp are shown in dark green.
Within a parking slot ${S}_{i,d,t}$, a truck may either wait to be charged, abandon charging, or charge directly using one of the available charger types $j\in\mathcal{J}$. In each parking time slot, the effective parking duration is less than half an hour, which restricts the maximum amount of energy that can be charged for this truck during this parking time window. Across the optimization horizon, the total energy of each truck scheduled to be charged must satisfy the truck’s trip-related energy consumption, considering the battery energy balance, battery SoC constraints, and charging power constraints. The traveling distance and energy consumption during parking slots ${S}_{i,d,t}$ is not considered. Furthermore, at any time $t$ on the day $d$, the number of chargers simultaneously used in zone $z$ in type $j$ cannot exceed the number of this type of chargers installed in the zone, $x_{z}^{j}$.
2.2 Set, index, and parameters
  • $\mathcal{J},j$: Set, index of EV charger types.
  • $\mathcal{Z},z$: Set, index of EV charging zones.
  • $\mathcal{I},i$: Set, index of electric trucks.
  • $\mathcal{D},d$: Set, index of days in the planning horizon.
  • $\mathcal{T},t$: Set, index of hours in day $d$.
  • $\mathcal{S}_{i,d,t}$: Set of discrete possible parking times for truck $i$ on day $d$ at the time $t$.
  • $E_{i}^{\text{bat}}$: Battery energy capacity (kWh) of truck $i$.
  • $\text{SoC}^{\text{max}}$: Maximum SoC of EV batteries.
  • $\text{SoC}^{\text{min}}$: Minimum SoC of EV batteries.
  • $P^{j}$: Power rating (kW) of charger type $j$.
  • $\eta^{j}$: Charging efficiency of charger type $j$.
  • $\rho_{i,d,t}$: Energy consumption usage (kWh) of truck $i$ at time $t$ on day $d$.
  • $C_{j}^{\text{install}}$: Installation cost for charger type $j$.
  • $P^{\text{low}}$: Penalty for low SoC inducing the range anxiety
  • $P^{\text{charging}}$: Penalty for charging time of trucks.
  • $\mu$: Mean of parking durations by truck $i$ at hour $t$ at zone $z$ across all the optimized days.
  • $\sigma$: Standard deviation of parking durations by truck $i$ at hour $t$ on day $d$ at zone $z$ across all the optimized days.
  • $\Delta t$: Duration (mins) in each parking time interval
2.3 Decision variables
  • $x_{z}^{j}\in\mathbb{Z}^{+}$: The number of chargers of type $j$ installed in zone $z$.
  • $w_{i,d,t}\in\mathbb{R}^{+}$: Accumulated waiting time for truck $i$ at charger $j$ on day $d$ at time $t$.
  • $b_{i,d,t}\in\mathbb{R}^{+}$: Battery SoC level (%) of truck $i$ at time $t$ on day $d$.
  • $p_{i,d,t}\in\mathbb{R}^{+}$: EV charging power (kWh) of truck $i$ at time $t$ on day $d$.
  • $v_{i,d,t}\in\mathbb{R}^{+}$: Battery SoC inadequacy below the low SoC threshold.
  • $T^{\text{max}}_{i,d,t}\in\mathbb{R}^{+}$: Maximum waiting time before truck $i$ abandons charging at time $t$ on day $d$.
  • $\delta_{i,d,t}^{30}\in\{0,1\}$: Indicator for the lower SoC bound of EV batteries inducing range anxiety.
  • $y_{i,d,t}^{j}\in\{0,1\}$: Binary variable indicating that truck $i$ is charged with charger type $j$ at time $t$ on day $d$ (1 for charging and 0 for not).
  • $a_{i,d,t}\in\{0,1\}$: Binary variable indicating that truck $i$ abandons charging at time $t$ on day $d$ (1 for abandonment and 0 for not).
2.4 Uncertainty variables
  • $pp_{i,j,z,t}\in\mathbb{R}^{+}$: Parking duration for truck $i$ at time $t$ in zone $z$ using charging type $j$.
2.5 Objective function
The objective function consists of two stages: the charging planning and scheduling stages. The planning stage considers multiple types of charging infrastructure, such as slow and fast chargers. In addition to power capacity, we also model the SoC charging range for each type of charger, which is detailed in the following section. The operational part includes penalties for low battery SoC, which induces range anxiety and affects charging time during scheduling.
[…]
where the low SoC penalty is calculated as the sum of the soft violation variable $v_{i,d,t}:=[0.3-b_{i,d,t}]^{+}$. This variable is greater than zero (i.e., $v_{i,d,t}>0$) if and only if the battery SoC level falls below 30 %. The charging time penalty is calculated as the sum of all the charging times across all zones and the time horizon to prevent low charger utilization.
2.6 Constraints for charging planning and scheduling
The following subsections detail constraints pertaining to charging planning and scheduling for the entire fleet of electric trucks. We also introduced three big-M constraints, and we discussed the range value of these $M$ in Appendix 7.
2.6.1 Truck charging constraints
We impose the following constraints to govern the truck charging behavior:
(Single charger usage) $\sum_{j\in\mathcal{J}}y_{i,d,t}^{j}\leq 1$ $\forall i,d,t$ (1)
(Charging only in parking windows) $y_{i,d,t}^{j}=0$ $\forall i,t\notin\mathcal{S}\quad\forall j$ (2)
(Low SoC indicator) $b_{i,d,t}-\frac{30}{100}\geq-M_{1}\cdot\delta_{i,d,t}^{30}+\epsilon$
$b_{i,d,t}-\frac{30}{100}\leq M_{1}\cdot(1-\delta_{i,d,t}^{30})$ $\forall i,d,t.$ (3)
Constraint (1) ensures each truck uses at most one charger at any time. Constraint (2) enforces charging only during the scheduled parking window. Constraint (3) defines the binary indicator $\delta_{i,d,t}^{30}$ for the low SoC condition, where $\delta_{i,d,t}^{30}=0$ if and only if the battery level exceeds 30%. Here, $M_{1}$ is a large constant, and the effective range of $M_{1}$ is presented in Appendix 7. The parameter $\epsilon$ is a small positive number, and it ensures that the indicator equals one only when the battery SoC is strictly greater than 30%, thereby guaranteeing a unique waiting time when the battery SoC is exactly 30%.
2.6.2 Abandonment and waiting time constraints
We model two coupled phenomena during each half–hour parking slot $t$ on day $d$: (i) abandonment—once a driver gives up waiting, no charging may occur in that slot; and (ii) waiting-time accumulation—while a truck remains parked in the same zone across consecutive admissible slots and is not charging, its waiting time increases by $\Delta t$; otherwise it resets to zero.
Abandonment
Let $a_{i,d,t}\in\{0,1\}$ indicate that truck $i$ abandons in slot $(d,t)$ and let $y^{j}_{i,d,t}\in\{0,1\}$ be the charger-type assignment. We enforce
(Once abandon, no charging) $a_{i,d,t}\;\leq\;1-\sum_{j\in\mathcal{J}}y^{j}_{i,d,t},$ $\forall i,d,t,$ (4)
so choosing abandonment excludes charging in that slot. During a continuous parked stretch (defined below), the abandonment variable is non-decreasing:
$a_{i,d^{\prime},t^{\prime}}\;\geq\;a_{i,d,t},\qquad\forall(d^{\prime},t^{\prime})\text{ successor of }(d,t).$ (5)
(Persistence over a parked stretch: once abandoned, a truck remains abandoned in subsequent slots within the same zone.)
Waiting time
Let $w_{i,d,t}\geq 0$ be the accumulated waiting time upon entering slot $(d,t)$. When two consecutive admissible slots for truck $i$ occur in the same zone, waiting evolves as
(Within-zone recursion) $w_{i,d,t+1}\;=\;w_{i,d,t}+\Delta t\!\left(1-\sum_{j\in\mathcal{J}}y^{j}_{i,d,t}\right),$ $\forall i,d,t\text{ with same-zone successor,}$ (6)
i.e., $w$ grows by $\Delta t$ only if the truck was not charging in the preceding slot. Otherwise, $w$ resets:
[…]
Rule (9) initializes $w$ at the start of any parked stretch and breaks carryover across zone changes; Equation (6) propagates $w$ across day boundaries when the zone is unchanged (e.g., from $(d,T_{d}{-}1)$ to (d+1,0)(d{+}1,0)). Two special cases for resetting the waiting time, the day-to-day carryover of waiting time in the same zone and zone-change reset of waiting time, are presented in Figs. 2 and 3, respectively. The full illustration of waiting time and abandonment is presented in Appendix 6.
Figure 2: Day-to-day carryover of waiting time in the same zone. For Truck 2 on Day 1, the first admissible slot initializes $w=0$. The next slot in the same zone accumulates to $w=\Delta t$, and carryover to Day 2 in the same zone yields $w=2\Delta t$.
Figure 3: Zone-change resets of waiting time. For Truck 2, waiting grows from $w=0$ to $w=\Delta t$ across two slots in Zone 1. Transition to Zone 2 resets $w=0$, after which accumulation resumes in subsequent uncharged slots.
2.6.3 Charging abandonment rules
At an industrial site, a truck abandons charging if its waiting time exceeds a maximum allowed waiting time $T^{\text{max}}_{i,d,t}$, as constraint (10). $M_{2}$ is a large constant to relax the waiting time constraint when $a_{i,d,t}$ equals to 1, in other words, when the trucks abandon charging. Notably, the maximum waiting time is a variable and it depends on the truck SoC due to the range anxiety and whether it is parked in special zones that allow for the overnight charging.
Based on experiences from the industrial partner, we formulate the following waiting and abandonment rules, considering range anxiety and overnight charging in the special zones. As constraint (11), if the SoC is greater than or equal to 30%, no waiting time is allowed as the driver has no range anxiety. If the SoC of a truck is below 30%, this will result in the driver’s range anxiety and would extend their waiting time to 0.5 hours before abandoning charging. Let the set $S^{Z}$ denote the parking time slots at any of the special zones for overnight charging. Constraint (12) sets the waiting time to be zero when the truck is outside the parking windows (i.e., no parking allowed) or in special zones (i.e., no waiting time limits). Constraint (13) regulates that the waiting time is always non-negative.
(Abandon after the maximum waiting time) $w_{i,d,t}\leq T^{\text{max}}_{i,d,t}+M_{2}\cdot a_{i,d,t}$ $\forall i,d,t\in\mathcal{S}$ (10)
(Maximum waiting time in non-special zones) $T^{\text{max}}_{i,d,t}=\Delta t\cdot\delta_{i,d,t}^{30}$ $\forall i,d,t\in\overline{\mathcal{S}^{Z}}$ (11)
(Outside parking windows or in special zones) $w_{i,d,t}=0$ $\forall i,d,t\in\mathcal{S}^{Z}\cup\overline{\mathcal{S}}$ (12)
(Non-negative constraint) $w_{i,d,t}\geq 0$ $\forall i,d,t\in\mathcal{S}$ (13)
2.6.4 Battery dynamics constraints
We impose the following constraints to model battery SoC evolution and charging power limits:
(Battery energy balance) $b_{i,d,t}E^{bat}_{i}=b_{i,d,t-1}E^{bat}_{i}+p^{ch}_{i,d,t}-\rho_{i,d,t},$ $\forall i,d,\forall t\geq t_{0}$ (14)
(The initial and final SoC are equal) $b_{i,d_{0},t_{0}}=b_{i,d_{e},t_{e}}$ $\forall i,d,t$ (15)
(Maximum charging power limit) $p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}\cdot P^{j}\cdot\Delta t\cdot y_{i,d,t}^{j}\cdot pp_{i,d,t},$ $\forall i,d,t\in\mathcal{S}$ (16)
(Effective charging time limits) $\frac{5}{30}\leq pp_{i,d,t}\leq 1,$ $\forall i,d,t\in\mathcal{S}$ (17)
(Battery SoC limits) $\text{SoC}^{\text{min}}\leq b_{i,d,t}\leq\text{SoC}^{\text{max}}$ $\quad\forall i,d,t$ (18)
Constraint (14) models EV battery dynamics considering charging power and energy consumption, and it should be noted that we do not use a binary variable to model battery discharging status, as it shall not happen during parking. Constraint (15) sets the initial SoC at $t_{0}$ and final SoC at time $t_{e}$ of the optimization horizon to be the same. Constraint (16) sets the maximum charging power limit considering the charging type $P^{j}$ and corresponding charging efficiency $\eta^{j}$. Constraint (17) regulates the charging time limit in each time window, which is greater than 5 mins and less than 30 minutes. Finally, constraint (18) restricts battery SoC within the maximum and minimum values.
2.6.5 Fast charger SoC restrictions
To prevent trucks from using fast chargers when their SoC exceeds $80\%$ and to ensure that fast chargers do not charge the battery SoC beyond $80\%$, the following constraints are enforced when charging. Constraint (19) ensures that if $b_{i,d,t}>80\%$, then $\delta_{i,d,t}^{80}=1$, and $y_{i,d,t}^{1}=0$, preventing the use of the fast charger. Also, if $b_{i,d,t}\leq 80\%$, then $\delta_{i,d,t}^{80}=0$, allowing the fast charger to be used. Constraint (20) ensures that when the fast charger is used, the battery level after charging ($b_{i,d^{\prime},t^{\prime}}$) is limited to at most 80%.
(SoC cap before fast charging) $b_{i,d,t}\leq\frac{80}{100}+M_{3}\cdot(1-y_{i,d,t}^{1})$ $\forall i,\;d,\;t\in\mathcal{S}$ (19)
(SoC cap during fast charging) $b_{i,d^{\prime},t^{\prime}}\leq\frac{80}{100}+M_{3}\cdot(1-y_{i,d,t}^{1})$ $\forall i,\;d^{\prime},\;t^{\prime}\in\mathcal{S}\;$ (20)
Here, $b_{i,d,t}$ is the SoC of truck $i$; $y_{i,d,t,1}$ is a binary variable equal to 1 if truck $i$ uses a fast charger at time $t$ on day $d$; $M_{3}$ is a large constant; and $\mathcal{S}$ is the set of time steps $t$ on day $d$ when truck $i$ is parked. The pair $(d^{\prime},t^{\prime})$ represents the next time step after $(d,t)$, defined as $(d,t+1)$ if t<T1t<T-1, or $(d+1,0)$ if t=T1t=T-1 and d<D1d<D-1, where $T$ is the number of time steps per day and $D$ is the number of days.
2.6.6 Decision-dependent uncertainty sets of parking duration
The effective charging duration $\hat{pp}$ is stochastic and influenced by the charging decision $y_{i,d,t}^{j}$, which is represented as binary variables. We constructed a box-shaped, decision-dependent uncertainty set for truck $i$ in zone $z$ at hour $t$. This is formulated as a robust linear constraint with affine decision dependence in terms of charging decisions, $y_{i,t,z}^{j}$.
$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}\cdot P^{j}\cdot\Delta t\cdot y_{i,d,t}^{j}\cdot\hat{pp}_{i,t,z},\quad\hat{pp}\in\Phi(y_{i,t,z}^{j})\quad\forall i,d,t\in\mathcal{S}$ (21)
The box-shaped uncertainty set $\Phi(y_{i,t,z}^{j})$ is defined as,
$\Phi=\left\{\hat{pp}\in\mathbb{R}^{i\times t}\ \Bigg|\ \frac{5}{30}\leq\hat{pp}\leq 1,\ \mu_{i,t,z}(y_{i,t,z}^{j})-\sigma_{i,t,z}(y_{i,t,z}^{j})\leq\hat{pp}\leq\mu_{i,t,z}(y_{i,t,z}^{j})+\sigma_{i,t,z}(y_{i,t,z}^{j})\right\}$ (22)
Where the bounds of the uncertainty set depend on the the minimum and maximum charging time, and calculated mean $\mu_{i,t,z}$ and standard deviation $\sigma_{i,t,z}$ associated with the charging decisions. The nominal moments are extracted from the historical data. Many charging trial data shows that fast charging could reduce the uncertainty of parking duration as the charging time becomes short. Therefore, we introduced a factor $\gamma$ to represent the reduction in uncertainty resulting from the fast charger usage.
$\mu_{i,t,z}(y_{i,t,z}^{j}):=\hat{\mu}_{i,t,z}\left(y_{i,t,z}^{0}+y_{i,t,z}^{1}\right)$ (23)
$\sigma_{i,t,z}(y_{i,t,z}^{j}):=\hat{\sigma}_{i,t,z}\left(\gamma^{0}\cdot y_{i,t,z}^{0}+\gamma^{1}\cdot y_{i,t,z}^{1}\right)$ (24)
Where $y^{0}$ and $y^{1}$ indicate charging decisions using slow or fast chargers, respectively. Notably, the decision-dependent RO is NP-complete, non-linear due to the multiplication of variables \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP. Here we proved that the proposed formulation is a special case that can be formulated as a robust linear constraint (33), and therefore the model can be formulated as a MILP problem. The proof is presented in Appendix 8.
$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}\cdot P^{j}\cdot\Delta t\cdot(\hat{\mu}_{i,t,z}\cdot y_{i,d,t}^{j}-\hat{\sigma}_{i,t,z}\cdot\gamma^{j}\cdot y_{i,t,z}^{j}),\quad\forall i,d,t\in\mathcal{S}$ (25)
2.6.7 Total number of chargers in each type and zone
The number of trucks charging with charger type $j$ cannot exceed the number of available chargers of type $j$ in zone $z$ at any time.
[…]
Where $\mathbbm{1}$ is a binary indicator to represent whether truck $i$ is at zone $z$ at time $t$ of day $d$.
3 Robust planning using the whole year dataset
The planned EV infrastructure must remain robust across the year-round fleet operations. We compare two approaches to solving based on the whole-year dataset: day selection (DS) and iterative planning and scheduling (IPS).
  • DS: We selected a certain number of representative days based on statistical criteria, such as the total fleet distance, the number of moving trucks, the total fleet energy consumption, and the total stopping time during a day. We also included days around the representative days to preserve the trip and battery cycle continuity.
  • IPS: We first decomposed the full-year dataset into monthly subsets, allowing several overlapping days between months. The algorithm then plans charger deployment for the entire year by iteratively integrating scheduling and monthly planning. It starts with an initial charger allocation derived from the dataset and processes each month sequentially. If the existing infrastructure can satisfy the charging demand of this month, only charging scheduling is performed. Otherwise, the algorithm jointly determines the charging schedule and the additional chargers required, updates the infrastructure, and carries the new configuration forward to subsequent months.
4 Fix-and-optimize heuristic based on high charging demand forecasts
Directly solving the proposed planning and scheduling model can be computationally demanding with a long optimization horizon. In this paper, we propose a fix-and-optimize heuristic designated for the scheduling problem based on high charging demand forecasts. The fix-and-optimize heuristic has been applied to timetable \@BBOPcite\@BAP\@BBNFonseca et al. (2024)\@BBCP and job-shape scheduling \@BBOPcite\@BAP\@BBNLi et al. (2025)\@BBCP problems, where part of the solution is fixed while the remaining variables are free to form subproblems. In this paper, as presented in Fig. 4, we identified the parking intervals with high charging demand at an early stage based on the samples from the uncertainty set and pre-assigned fast-charging schedules, thereby accelerating the overall optimization process.
Figure 4: The fix-and-optimize heuristic where fast-charging scheduling is pre-assigned and fixed based on the forecast high charging demands
As shown in Algorithm 1, we first initialize the master problem using the starting values $pp_{i,d,t}$ to solve the model. Next, we sample the stochastic parking duration $\hat{pp}_{i,d,t}$ from the uncertainty set $\Phi(\mu,\epsilon)$ to replace the initial values and compute the upper bounds of the charging power based on these samples. We then compare the required charging power against the updated upper bounds to identify any violations. A violation occurs if the required charging power exceeds the upper bound by more than a threshold $\kappa^{v}$ and the parking window does not belong to a special zone. The threshold $\kappa^{v}$ is initialized at two and decays exponentially across iterations. When a violation is detected, a fast-charging scheduling constraint $y_{i,1,d,t}=1$ is added for that parking window. These constraints are subsequently incorporated into the master problem, after which the model is reset and resolved from the first step.
Input: Parking duration samples $\hat{pp}_{i,d,t}$ drawn from the uncertainty set $\Phi(\mu,\epsilon)$
Output: Charger installation decisions $x_{z}^{j}$
repeat
   Solve the proposed optimization problem with deterministic $pp_{i,d,t}$ as the initial input (master problem); $\triangleright$ Warm start if the solution of the master problem available;
   Obtain scheduling outcomes $y_{i,j,d,t}$ and charging profiles $p^{ch}_{i,j,d,t}$;
  Sample parking durations $\hat{pp}_{i,d,t}$ from the uncertainty set $\Phi(\mu,\epsilon)$;
 for $i\in\mathcal{I},\;t\in\mathcal{T},\;d\in\mathcal{D},\;j\in\mathcal{J}$ do
      Compute the new upper bound of charging power:
$p^{ch^{\prime}}_{i,j,d,t}\leftarrow\sum_{j\in\mathcal{J}}\Delta t\cdot P^{j}\cdot y_{i,j,d,t}\cdot\hat{pp}_{i,d,t}\cdot\eta^{j}\;$
    if $\big(p^{ch}_{i,0,d,t}\geq p^{ch^{\prime}}_{i,0,d,t}+\kappa^{v}\big)\land(\forall i,d,t\in\overline{\mathcal{S}^{Z}})$ or $p^{ch}_{i,1,d,t}\geq p^{ch^{\prime}}_{i,1,d,t}$ then
         Impose the fast-charging constraint $y_{i,1,d,t}=1$ in the master problem;
until The maximum iteration limit is reached;
Post-process the solution by removing fast charging results in zones where the number of fast chargers exceeds two;
Re-solve the optimization problem using the sampled values $\hat{pp}_{i,d,t}$ and incorporating the charger installation constraints derived from the processed solution;
Algorithm 1 Fix-and-optimize heuristic under effective parking duration uncertainty
This iterative process continues until the maximum number of iterations is reached. Afterward, we post-processed the optimization results by removing zones with more than two fast chargers. The remaining fast charger installations were treated as hard constraints, while the solutions for slow chargers were used as lower bounds. Ultimately, the installation plan for all zones and charger types was determined.
5 Application
We applied our model to an open-pit mining site with 203 light-duty trucks in operation, which are planned to be replaced by battery EVs. Seven special-use trucks, such as ambulances, that will not be electrified, along with another seven trucks dispatched fewer than 10 days, are excluded from the study. The remaining 189 trucks will be replaced with electric trucks equipped with 100 kWh batteries. We consider the following two types of chargers, as listed in Table 1. The capital cost of fast chargers includes the installation cost. We set the penalty of low SoC and charging time to be one in the objective function. The charging efficiency could vary with the charging type, outdoor temperature, and charging power \@BBOPcite\@BAP\@BBNSevdari et al. (2023)\@BBCP. This paper uses a fixed value for charging efficiency. Overall, fast charging tends to be more efficient than slow charging, as it uses a high-efficiency rectifier located outside the truck \@BBOPcite\@BAP\@BBNArena et al. (2024)\@BBCP. We therefore set the charging efficiency of slow and fast charging to be 90% and 95%.
Specifications Slow charger Fast charger
Capital costs (\$/charger) 1,500 \@BBOPcite\@BAP\@BBNFord (2025)\@BBCP 38,000 \@BBOPcite\@BAP\@BBNPrimecomtech (n.d.)\@BBCP
Maximum charging rates (kW) 11.2 150
SoC ranges 10–100% 10–80%
Charging efficiency 90% \@BBOPcite\@BAP\@BBNSevdari et al. (2023)\@BBCP 95% \@BBOPcite\@BAP\@BBNArena et al. (2024)\@BBCP
Table 1: Technical parameters of two types of EV chargers
The charging planning process begins by identifying potential charging sites based on the stop frequency of all trucks (Section 5.1). Next, the parking window, effective parking duration, and energy consumption for each truck are extracted (Section 5.2). Using this information, the DS and IPS methods are applied to plan across the full-year horizon. If the RO formulation is used, a decision-dependent uncertainty set is constructed from the historical samples (Section 5.3). High-consumption vehicles with infeasible trips are then filtered out (Section 5.4). Finally, chargers are planned through the proposed model in the deterministic or stochastic formulation (Section 5.5).
5.1 Identifying potential EV charging points
We used GPS records of 189 light-duty trucks at the mining site over one year. Their GPS data records include coordinates (i.e., longitude and latitude) and speed at each time step. We then grid-divided the entire mining site into 100-meter $\times$ 100 100-meter square regions and counted the stop frequency in each square. We assumed that a truck stops if it moves less than 50 meters within a square region for at least five minutes. Comparing the number of stop frequencies in each square, we identified eight EV charging points with the highest number of stops throughout the year. Fig. 5 shows eight EV charging points identified, and the dashed gray lines indicate the roads connecting them. Fig. 6 shows the total stopping time of the entire fleet ($\times$ 1,000 hours), including both short parking durations (<< 4 hours) and long parking durations ($\geq$ 4 hours), along with the annual number of trucks stopping at each zone. The detailed statistics of eight charging zones are listed in Table 6 in Appendix 9.1. The zone 1 (roadhouse) and zone 4 (equipment maintenance) have a large part of long duration time greater than 4 hours, which means that most of the overnight parking happens at zones 1 and 4.
Figure 5: Identified eight charging zones and roads connected to eight charging zones (indicated as gray dashed lines)
Figure 6: The total stopping time of the entire fleet ($\times$ 1,000 hours), including both short parking durations (<< 4 hours) and long parking durations ($\geq$ 4 hours), along with the annual number of trucks stopping at each of eight zones
5.2 Summary of input datasets
We extracted three inputs for each truck in each half-hour interval based on their GPS records: (i) the travel distance of each truck in miles, (ii) the parking time windows as the input of the set $\mathcal{S}_{i,d,t}$, and (iii) the effective parking duration in each time window in minutes as the input of $pp_{i,d,t}$. The travel distance is then converted to energy consumption for each half-hour parking time window based on the second-order model of the temperature-dependent fuel economy for U.S. trucks \@BBOPcite\@BAP\@BBNGoodall & Robartes (2024)\@BBCP, as the input of $\rho_{i,d,t}$. The detailed procedures to extract all three kinds of inputs are illustrated in Appendix 9.3.
Figure 7: (a) The histogram of average daily distance per truck (miles) and (b) the histogram of average daily energy consumption per truck (kWh)
Figs. 7 (a) and (b) illustrate the distributions of average daily distance traveled per truck and the corresponding average daily energy consumption. On average, each truck traveled 73.9 miles per day. With a fuel economy of roughly 3–3.4 miles/kWh, this translates to a daily energy use of 22.4–24.6 kWh per truck, or about 30% of the battery’s capacity. This implies that a full battery cycle would typically last 2–3 days. It is important to note that travel distance and energy consumption within charging zones are excluded from the model. We expect, however, that this additional consumption could be offset with minor operational adjustments, such as modifying parking durations.
5.3 Samples for constructing robust uncertainty sets
To construct a robust uncertainty set for heterogeneous effective parking durations, we first collected all parking durations and parking window data across the year. Fig. 8 (a) illustrates the total time of actual parking duration versus the the total time of the full parking window per truck per day at eight zones. In each zone, a portion of the parking window cannot be used for charging, for example, when the truck is moving to find a parking spot. Fig. 8 (b) presents the distribution of parking durations across eight zones, showing that zones 6, 7, and 8 exhibit greater variability compared to the others.
Figure 8: (a) Actual parking duration versus full parking windows duration per day per truck and (b) box plots of actual parking duration distributions of eight zones
Figure 9: (a) Mean and (b) variance of effective parking durations in minutes of 189 trucks by hour of the day in the charging zone 1
We found that the actual parking duration at the mining site is strongly related to trucks, hour of the day, and charging zones, but has almost no correlation with day of the week or day of the month. This is because the mining site operates continuously with the same pattern throughout the year. We then calculate the mean and standard deviation of parking durations pertaining to trucks, hours of the day, and zones. Fig. 9 shows the mean and variance of effective parking duration in minutes by hour of the day in the charging zone 1. Most trucks operating around the mining site have zero means and variances in their actual parking duration during the day.
5.4 Set up
We prepared three inputs based on the dataset of the year 2023, for the DS and IPS approaches, respectively. For the DS approach, we defined 10 statistical criteria, including total fleet distance, number of active trucks, and total fleet energy consumption. Based on these criteria, we ranked all 365 days in a year and selected the top-ranking days. We experimented with different numbers of selected days, ranging from 10 to 40, and observed that charger installation results stabilized once the selection reached 30 days. Accordingly, we chose 39 days, including days adjacent to the top 10 ranked days, as some vehicles charge every two or three days, and representative sub-periods need to extend beyond battery cycles. The detailed selection criteria and procedures are provided in Appendix 9.2. For the IPS approach, we divided the full-year dataset into monthly subsets, incorporating overlapping days between consecutive months to maintain trip continuity. This procedure produced 12 monthly datasets, each spanning 34 days.
We conduct a preliminary simulation before optimization, in which each EV uses straightforward fast charging throughout the optimization horizon. In each simulation, a truck starts with a random SoC and is then fully charged at each parking site. At the end of the optimization horizon, we track whether they could restore their original SoC. If the original SoC cannot be recovered, the truck is identified as a high consumption vehicle (HCV), which cannot be feasibly electrified if it sticks to the current operational scheduling. This also includes vehicles not passing the eight charging zones. The optimization is subsequently performed exclusively on the rest of the trucks, and we recorded the number of HCVs excluded from the optimization model in each simulation. All experiments were solved using the Gurobi commercial solver. We set the accepted optimal gap for this MILP to be 1%, less than the capital cost of one slow charger. All problems were solved on a computer equipped with an 8-core CPU running at 3.4 GHz and 160 GB of RAM.
5.5 Result evaluation
5.5.1 The impacts of abandonment, range anxiety, and effective parking duration
We first evaluate four deterministic cases to analyze the impacts of charging abandonment, range anxiety, and parking duration restrictions. Using the DS method, we draw on 39 selected representative days for the input datasets. The total energy consumption of the entire vehicle fleet for the 39-day optimization horizon is 83,050 kWh, which is around 11.33 kWh per day per truck. The total parking time of the entire vehicle fleet is 124,639 hours, which is around 17 hours per day per truck.
  • Case 1 - Benchmark: Trucks are prohibited from waiting in any zone under normal circumstances. However, when a vehicle SoC drops to below 30%, it is permitted to pause for 30 minutes to recharge. Two designated zones (i.e., zone 1 and zone 4) accommodate the overnight charging with unlimited waiting time. The model incorporates the specific parking duration for each parking window (i.e., $\frac{5}{30}\leq pp\leq 1$).
  • Case 2 - Full parking time: This case maintains the same conditions as the benchmark case, except that it has no specific parking duration constraints for each parking window (i.e., $pp=1$).
  • Case 3 - No overnight charging: This case maintains the same conditions as the benchmark case, except that it has no special zones for the overnight charging.
  • Case 4 - No range anxiety: This case maintains the same conditions as the benchmark case, except that it has no related constraints or penalty related to the range anxiety, in order words, there is no penalty for low SoC (i.e., $P^{low}=1$). Trucks are prohibited from waiting in any zone under normal circumstances, regardless of their SoC. In other words, the maximum waiting time is set to zero (i.e., $T^{max}_{i,d,t}=0$).
Table 2 summarizes optimization results in four deterministic cases under different assumptions about abandonment, range anxiety, and effective parking duration. The benchmark case (case 1) has a mixed type of chargers, which are 72 slow chargers and 7 fast chargers. Compared to the benchmark case, case 2 shows that a less number of fast chargers are needed if the model does not consider effective charging time in each parking time slot. In case 3, an additional number of 20 slow chargers are installed if there is no special zone allowing for overnight charging. In case 4, if the model does not consider range anxiety, seven additional slow chargers need to be installed to accommodate more frequent abandonment without waiting time.
Instances
# of
HCVs
# of slow /
fast chargers
Installation
cost (\$)
Scheduling
penalty (\$)
Gap (%) Time (h)
Case 1 1 72/6 336,000 38,468 0.084 2.98
Case 2 1 74/3 187,000 39,622 0.116 57.32
Case 3 1 92/6 366,000 85,200 0.042 15.77
Case 4 1 79/6 346,500 38,077 0.079 2.94
Table 2: Summary of optimization results in four deterministic cases: benchmark, no parking time restriction, no overnight charging, and no range anxiety
Figure 10: Zone-wise differences in the number of slow and fast chargers of cases 2-4 compared with the benchmark
Fig. 10 illustrates the differences in the numbers of fast and slow chargers in cases 2–4, compared with the benchmark. Across four cases, the number of installed chargers is proportional to the total duration of vehicle stops at each charging zone, as presented in Fig. 8 (a), and zones 4 and 5 have the highest number of chargers. In case 1, setting the special zones does not relax the requirement of the fast charger installation in these zones, as in zones 1 and 4. In case 3, without considering the overnight charging, zones 1 and 4 will have a high number of slow chargers. In case 4, without considering the range anxiety, zones 5-7 will have a higher number of slow chargers, meaning vehicles passing through these three zones have a relatively low SoC than other zones.
Instances
Total slow
charging
hours
Total fast
charging
hours
Total
abandonment
hours
Avg. charging
hours
(truck per day)
Avg. charging
power
Case 1 3677.0 1352.5 26371.0 0.69 16.51
Case 2 5779.0 555.5 26430.5 0.86 13.11
Case 3 3493.5 1418.5 69353.0 0.67 16.90
Case 4 3116.5 1248.5 26947.0 0.60 19.02
Table 3: Summary of the scheduled charging hours, abandonment hours, and charging power across four deterministic cases
Figure 11: The total number of (a) charging hours and (b) abandonment hours of the entire truck fleet in a day across four instances
Figure 12: The total number of (a) charging hours and (b) abandonment hours of the truck fleet in different hours of a day in four deterministic cases
Figure 13: (a) Average SoC of each truck and (b) the overall battery SoC for the entire fleet in four deterministic cases
We calculate the charging power, charging hours, and abandonment hours across four deterministic cases, summarized in Table 3. The detailed scheduling solution of four cases are presented in Appendix 10. Case 2, which assumes full parking duration, results in the highest total charging hours but the lowest average charging power due to having the fewest fast chargers. Case 3, which excludes overnight charging, produces the greatest number of abandonment hours. Case 4, which removes range anxiety, leads to the fewest charging hours and the highest average charging power.
Figs. 11 (a) and (b) show the zone-wise comparison of charging and abandonment time, respectively. In case 2, with the full parking duration, zones 6 and 8 have a higher charging time. In case 3, without the overnight charging, zones 1 and 4 have a significantly high number of abandonment hours. Figs. 12 (a) and (b) show the total number of charging and abandonment hours of the total truck fleet in different hours of a day in four cases, respectively. Across four instances, the lowest number of charging hours appears at 12:00, and the highest number of charging hours appears at 22:00. If the model does not assign two special zones for overnight charging, the number of abandonment hours will increase by more than a factor of 2.
Fig. 13 (a) illustrates the average SoC of all trucks over 39 days, ranked from the highest to lowest, and Fig. 13 (b) presents the overall battery SoC (maximum, minimum, and average) of the entire fleet in four cases. Among four cases, case 4 (no range anxiety) yields the lowest average SoC ($28.4\%$), whereas case 2 with full parking duration achieves the highest average SoC ($45.3\%$). When the effective charging time constraint is applied, in other words, comparing case 2 to case 1 benchmark, part of the fleet experiences a sharp drop in the average SoC to very low levels ($\leq 20\%$), due to the limited charging duration in the parking windows.
5.5.2 Robust infrastructure planning based on the whole year dataset
[…]
To ensure robust infrastructure planning throughout the year, we optimize the charging infrastructure using two solving approaches: the DS and IPS methods, as described in Section 3. For the DS method, we selected a 39-day horizon and imposed a maximum time limit of 40,000 seconds to solve each instance, recording the gap whenever the solver failed to converge within this limit. For the IPS method, we set a 6-hour time limit for each monthly instance. In addition, we applied the proposed fix-and-optimize heuristic to the 39-day instance under the DS framework, with the initial master problem aligned with the deterministic benchmark case.
Instances
($\sigma$, $\gamma$)
Solving
methods
# of
HCVs
# of slow/fast
chargers
Installation
costs (\$)
Scheduling
penalty (\$)
Gap (%)
Time (h)
(0, 1) DS 19 68/6 330,000 35,678 9.06 11.1
DS - h 19 68/6 330,000 35,679 0.07 2.40
IPS 48 77/5 305,500 33,466 0.09 11.3
(1, 1) DS 19 68/6 330,000 35,658 9.04 11.1
DS - h 19 69/6 331,500 36,065 0.05 2.15
IPS 48 75/7 378,500 33,331 0.05 12.0
(2, 1) DS 19 68/6 330,000 35,736 9.02 11.1
DS - h 19 69/6 331,500 35,863 0.05 5.20
IPS 48 76/8 418,000 33,402 0.05 23.8
(2, 0.5) DS 19 69/5 293,500 35,931 0.06 7.42
DS - h 19 68/6 330,000 35,715 0.06 3.29
IPS 48 74/8 415,000 22,541 0.03 29.8
Table 4: Summary of optimization results across four instances using DS, DS with heuristic (DS-h), and IPS methods. For the IPS method, the reported gap is the optimal gap of the final instance in the iterative planning
Figure 14: The total number of chargers in each month using the IPS method in four instances (in colored lines) versus the total number of trucks parking and charged (in gray bars) in all eight zones
[…]
Figure 15: Zone-wise differences in the number of slow and fast chargers in four stochastic cases compared to the benchmark
When comparing the three solution methods, the IPS method yields the highest number of chargers across all four instances using the full-year datasets. The DS and DS-h methods have an even lower number of chargers compared with the deterministic benchmark case, since an additional 18 trucks are excluded as HCVs due to the short parking duration. Fig. 14 shows the total number of chargers (i.e., the sum of fast and slow chargers) determined by the IPS procedure from monthly subproblems over 12 months in the four instances. From January to December, the number of trucks parking and charged in any of the eight zones rises from 156 to 182, leading to a steady increase in the required charger installations. As the parking duration uncertainty grows, the total number of EV chargers—particularly fast chargers—also increases. In the fourth instance, a larger share of fast chargers is installed compared to the third instance, since fast chargers are less affected by uncertainty than slow chargers. Across four stochastic instances, the only difference between the DS and DS-h results lies in the installation or placement of a single fast charger. This indicates that the assignment of specific fast chargers can hinder the convergence of the DS method. In contrast, the proposed heuristic prioritizes placing the fast charger, allowing the model to achieve the optimal solution more efficiently. The detailed solution process using the DS-h method is presented in Appendix 11.
Fig. 15 shows the zone-wise differences in the number of slow and fast chargers when comparing four stochastic instances to the benchmark case (i.e., case 1 in the deterministic formulation). In general, zones 1, 2, 4, and 7 have an increase in the number of slow chargers, while zone 6 has a reduction. It is because that zone 6 has a quite high variance in the parking duration. The model optimizes more chargers and schedules more trucks to be charged at zones 1 or 2, which are connected to zone 6, in four stochastic cases.
5.5.3 The impacts of uncertain parking duration
To evaluate the impact of uncertain parking duration on the charger planning and scheduling, we first simulated the monthly schedule based on the charger planning results for the case where $\sigma=2$ and $\gamma=0$, using the DS-h and IPS methods, respectively. We found that the charging installation obtained from the DS-h method meets the requirements for only six months, where the number of trucks dispatched is less, whereas those obtained from the IPS method satisfy the requirements for all 12 months. This discrepancy arises from the steadily increasing trend of truck parking and charging activities throughout the year, which the day-selection approach fails to capture adequately during peak months.
Instances
($\sigma$, $\gamma$)
Avg. slow
charging
hours
Avg. fast
charging
hours
Avg.
abandonment
hours
Avg. charging
hours
(truck per day)
Avg. charging
power (kW)
(0, 1) 3757.6 1280.7 20212.9 0.83 8.78
(1, 1) 2166.9 1249.9 20229.9 0.56 12.95
(2, 1) 2126.3 1357.5 20155.1 0.57 12.70
(2, 0.5) 515.1 1033.0 20191.6 0.25 28.59
Table 5: Summary of the average scheduled charging hours, abandonment hours, and charging power in a year across four instances
Figure 16: Boxplots of the monthly slow and fast charging hours across four stochastic instances versus the average of charging hours by month
We then evaluated four stochastic instances using the IPS method by simulating the charging scheduling in 12 months and calculated the average value of the total charging and abandonment hours by month. Fig. 16 shows the box-plots of the total number of slow and fast charging hours by months in four stochastic instances using the IPS method, and Table 5 summarizes the average scheduled charging hours, abandonment hours, and charging power by months. As the uncertainty increases where the standard deviation of the parking duration $\sigma$ increases from 0 to 2, the fast charger usage rises while the slow charger usage declines, since the shorter parking durations incentivize drivers to rely more on the fast charging. As the fast chargers are more frequently used, the average charging power increases from 8.78 kW when the standard deviation equals 0, to 12.95 kW and 12.70 kW when the standard deviation equals 1 and 2, respectively. The average abandonment hour keeps the same level across four stochastic instances.
For the second and third instances, where the standard deviations are 1 and 2, respectively, the third instance exhibits longer total charging hours and lower charging rates, even though fast-charging hours are greater than in the second instance. This is because the increased uncertainty in parking duration prolongs overall charging time, offsetting the benefit of more frequent fast charging. In contrast, in the fourth instance—where fast chargers effectively mitigate parking duration uncertainty—the total charging time decreases significantly, with trucks relying on fast chargers for most of the period. The average monthly slow-charging time is only 515.1 hours.
5.6 Conclusion and discussion
We developed a two-stage robust charging infrastructure planning model for the electrification of light-duty trucks at industrial sites with mixed-type chargers with different charging power. The first stage plans the charging infrastructure based on the selected charging locations and charging types. The second stage optimizes the charging schedules considering different waiting times, charging abandonment, and range anxiety. A case study is conducted on an open-pit mining site with eight potential charging zones. We used three input datasets extracted from the whole-year GPS data from around 200 trucks: (i) parking indications at each half-hour interval, (ii) actual parking durations of each parking window, and (iii) half-hour energy consumption of all trucks.
5.6.1 The impacts of overnight charging and range anxiety on charging behaviors
Among eight zones, we assume that two designated zones allow overnight parking; otherwise, trucks will forgo charging if no chargers are available, unless their battery level falls below 30%. The battery SoC status below 30% is assumed to trigger the range anxiety, resulting in an additional half-hour waiting time. The benchmark case based on the selected 39 days yields 72 slow chargers and 6 fast chargers. We then examine three cases where overnight charging, range anxiety, and limited parking duration in each parking window are not considered, respectively. Extending the parking duration to the full window reduces the number of fast chargers to 3. Removing the designated overnight zones leads to a dramatic increase in abandonment hours by 1.63 times and an additional 20 slow chargers installed in total. No range anxiety requires seven more slow chargers to be installed. In addition, we found that the charging hours only occupy around 4% of the total parking time, meaning the limited waiting time and consequential abandonment are the main reasons that lead to a number of slow chargers at industrial sites.
5.6.2 Robust charging infrastructure planning considering uncertain parking duration
To utilize the full-year dataset, we proposed two methods—the DS and IPS approaches—for charging infrastructure planning. The uncertain parking duration within each half-hour parking window is represented using a robust uncertainty set, whose bounds depend on the charging decisions. The IPS method ensures robust scheduling performance throughout the year by planning for a larger number of chargers, whereas the DS method fails in several months when the charging demand peaks.
When the uncertainty of charging duration increases (i.e., the lower bound of the uncertainty set decreases), the total number of fast-charging hours rises, while the total number of slow-charging hours declines. Consequently, the overall average charging power increases under the mixed-type charging infrastructure. However, further reducing the lower bound of the robust set for parking duration to the minimum level (i.e., 5 minutes) leads to longer total charging hours and a decrease in average charging power. If the model incorporates the mitigating effect of fast charging on parking duration uncertainty—observed in many real-world studies—a smaller number of slow chargers would be planned and utilized.
5.6.3 Heuristic methods to assist the robust scheduling with a long optimization horizon
The proposed two-stage planning and scheduling framework is formulated as a large-scale MILP model. For a 39-day scheduling instance with 189 trucks, the model contains over 1.5 million binary variables without pre-solving, making it computationally challenging. To address this, we introduce a fix-and-optimize heuristic for the charging scheduling under uncertainty. This is to forecast the high charging demand due to reduced parking duration and to pre-assign fast charging actions in the scheduling. This reduces the computation time by approximately 2-5 times for 39-day instances and yields the optimality gap of the problem below 0.1%.
Finally, several promising avenues remain for future research. For example, future work could incorporate additional charging options, such as multiple types of fast chargers with rated powers exceeding 150 kW, to further reduce the number of required HCVs. Moreover, charging behavior modeling could be improved by integrating empirical data from existing electrified fleets to account for factors like variations in range anxiety at different times of day.
5.7 Acknowledgment
We gratefully acknowledge the sponsor company for providing GPS data from the operating trucks at the mining site. We also thank Maria Isabel Castrillo Viguera for her valuable contributions in processing the GPS data and preparing the input datasets.

### v2

@@
 […]
-Decision-dependent Robust Charging Infrastructure Planning for Light-duty Truck Electrification at Industrial Sites: Scheduling and Abandonment
+Decision-Dependent Robust Charging Infrastructure Planning and Scheduling for Light-Duty Truck Electrification at Industrial Sites
 […]
@@
 […]
-Many industrial sites rely on diesel-powered light-duty trucks to transport workers and small-scale facilities, which has resulted in a significant amount of greenhouse emissions (GHGs).
-To address this, we developed a two-stage robust charging infrastructure planning model for electrifying light-duty trucks at industrial sites.
-The model is formulated as a mixed-integer linear programming (MILP) that optimizes the charging infrastructure, selected from multiple charger types and potential locations, and determines opportunity charging schedules for each truck based on the chosen infrastructure.
-Given the strict stopping points and schedules at industrial sites, we introduced a scheduling-with-abandonment problem, where trucks forgo charging if their waiting times exceed a maximum threshold.
-We also further incorporated the impacts of overnight charging and range anxiety on drivers’ waiting and abandonment behaviors.
-To represent stochastic, heterogeneous parking durations of trucks, we constructed a decision-dependent robust uncertainty set in which parking time variability flexibly depends on drivers’ charging choices.
-We applied the model in a case study of an open-pit mining site, which plans charger installations in eight zones and schedules a fleet of around 200 trucks.
-By decomposing the problem into monthly subproblems and using heuristic approaches, for the whole-year dataset, the model achieves an optimality gap of less than 0.1% within a reasonable computation time under diverse uncertainty scenarios.
+Many industrial sites and digital logistics platforms rely on diesel-powered light-duty trucks to transport workers and small-scale facilities, which results in a significant amount of greenhouse gas emissions (GHGs).
+To address this, we develop a robust model for planning charging infrastructure to electrify light-duty trucks at industrial sites.
+The model is formulated as a mixed-integer linear program (MILP) that optimizes the charging infrastructure selection (across multiple charger types and locations) and determines charging schedules for each truck based on the selected infrastructure.
+Given the strict stop times and schedules at industrial sites, we introduce a scheduling-with-abandonment problem in which trucks forgo charging if their waiting time exceeds a maximum threshold.
+We further incorporate the impacts of overnight charging and range anxiety on drivers’ waiting and abandonment behaviors.
+To model stochastic, heterogeneous parking durations, we classified trucks using machine learning (ML) methods based on contextual and time-location features.
+We then constructed decision-dependent, feature-driven robust uncertainty sets in which parking-time variability varies flexibly with drivers’ charging choices.
+These feature-driven sets are applied to two robust optimization formulations with decision-dependent uncertainty (RO-DDU), resulting in distinct outcomes and managerial implications.
+We conduct a case study at an open-pit mining site to plan charger installations across eight charging zones, serving approximately 200 trucks.
+By decomposing the problem into a short rolling horizon or using a heuristic approach for the full-year or representative-day dataset, the model achieves an optimality gap of less than 0.1% under diverse uncertainty scenarios.
+[…]
+Charging infrastructure planning; Decision-dependent uncertainty; ML classification; Contextual (Stochastic) optimization, Scheduling
 1 Introduction
 1.1 Background
-Diesel-powered light-duty trucks are used for transporting and dispatching goods in urban environments and industrial sites.
-However, their usage could produce a significant amount of GHG emissions.
-In 2022, light-duty trucks were the most significant GHG emission source of transportation, contributing to 58% of the total emission of the transportation sector in the U.S.
-\@BBOPcite\@BAP\@BBNU.S.
+Many industrial sites, such as the mining and metals sectors, and digital logistics platforms, rely on diesel-powered light-duty trucks to transport and dispatch goods.
+Their use has produced significant GHG emissions.
+In 2022, light-duty trucks were the largest GHG emission source in the U.S.
+transportation sector, contributing 58% of total emissions \@BBOPcite\@BAP\@BBNU.S.
 EPA (2024)\@BBCP.
-Battery electric vehicles (BEVs) play a central role in decarbonizing the transportation sector.
-Its associated charging infrastructure is currently categorized into two types.
-The first type is the level 1 and level 2 AC charging (hereafter referred to as ‘slow charging’), which generally requires between four hours and a full day to charge a vehicle battery from empty to 80% state of charge (SoC).
-The second type is the DC fast charging (hereafter referred to as ‘fast charging’), available at dedicated stations, which requires a relatively complicated installation process.
-The fast chargers can recharge an empty vehicle battery to 80% SoC in an hour \@BBOPcite\@BAP\@BBNUS Department of Transportation (2025)\@BBCP.
-Fast charging infrastructure also entails considerable capital investments and installation expenditures, primarily due to the power electronics required.
-Furthermore, its operational constraints differ from those of slow charging, as charging beyond 80% SoC is strongly discouraged for safety considerations \@BBOPcite\@BAP\@BBNMussa et al.
+Battery-electric vehicles (BEVs) are a key solution for decarbonization.
+Currently, the BEV charging infrastructure can be categorized into two types.
+The first type is Level 1 and Level 2 AC charging (hereafter referred to as “slow charging”), which generally requires between four hours and a full day to charge a vehicle battery from empty to 80% state of charge (SoC).
+The second type is the direct current fast charging (hereafter referred to as “fast charging”), which has been deployed at certain public stations.
+Fast chargers can recharge an empty vehicle battery to 80% SoC in about an hour \@BBOPcite\@BAP\@BBNUS Department of Transportation (2025)\@BBCP.
+However, fast-charging infrastructure entails considerable capital investment and installation expenditures, primarily due to the required power-electronics installations.
+Fast charging beyond 80% SoC is strongly discouraged for safety reasons \@BBOPcite\@BAP\@BBNMussa et al.
 (2017)\@BBCP.
-Another recharging option is the battery swapping station.
-The battery swapping station enables EVs to replace depleted batteries with fully charged ones swiftly \@BBOPcite\@BAP\@BBNMak et al.
+Another fast recharging option is battery swapping.
+Battery swapping enables EVs to replace depleted batteries with fully charged ones swiftly \@BBOPcite\@BAP\@BBNMak et al.
 (2013)\@BBCP.
-However, due to the safety risks associated with battery swapping operations, the feasibility of deploying battery swapping stations at industrial sites should be assessed on a case-by-case basis.
-In this paper, we examine the electrification of light-duty trucks at industrial sites for decarbonization.
-The original diesel-powered trucks are primarily employed for transporting workers and small-scale facilities.
-The strict operational schedules at industrial sites leave little room for trucks to change any designated parking locations or timetables.
-As a result, the scheduled trucks are expected to abandon charging if the anticipated waiting time for getting a charger surpasses a specified threshold.
-This is formulated as a ‘scheduling-with-abandonment’ problem.
-We conducted a case study of the open-pit mining site with 200 diesel-powered operational trucks.
-Without modifying any of their GPS data, we extracted their parking and driving patterns based on potential charging points and estimated their energy consumption at half-hour resolutions.
-Then, we developed a two-stage charging infrastructure planning and scheduling model for the entire truck fleet.
-The planned EV charging infrastructure satisfies the charging needs of each truck.
-That is meeting the energy consumption of each truck at any time, given the limited parking durations and battery energy capacity.
+However, compared with adding more trucks to the fleet, deploying battery swapping at industrial sites may yield limited benefits due to operational risks, including potential battery damage during swapping.
+In this paper, we develop an integrated planning and scheduling model for charging infrastructure with mixed-type BEV chargers, leveraging contextual stochastic optimization and ML techniques.
+We comprehensively incorporate heterogeneous BEV driving behaviors in the scheduling, including charging abandonment, overnight charging, and range anxiety.
+To capture variability in parking duration and its dependence on charging decisions, we construct a feature-based, decision-dependent uncertainty set using ML-based clusters and apply it to two RO-DDU formulations with different assumptions and computational costs.
+We conduct a case study at an open-pit mining site with 200 operational diesel-powered trucks.
+Without modifying the GPS data, we extract parking and driving patterns relative to potential charging locations, estimate energy consumption, and efficiently compute the charging infrastructure plan and schedules.
 1.2 Related Work
 1.2.1 Mixed-type Charging Infrastructure Planning
-Existing studies on charging infrastructure planning primarily focus on determining the location, number, and type of EV chargers, as well as the charging schedule using the selected installation.
-For electric buses and trucks at industrial sites, the charging process primarily involves opportunity or overnight charging, meaning they are only charged during their layovers and overnight parking \@BBOPcite\@BAP\@BBNAbdelwahed et al.
+Existing studies on charging infrastructure planning primarily focus on determining locations, quantities, and types of BEV chargers, as well as the associated charging schedules given selected installations.
+For electric buses and trucks at industrial sites, their charging operations typically rely on the opportunity or overnight charging, meaning that vehicles are charged during layovers and overnight parking periods \@BBOPcite\@BAP\@BBNAbdelwahed et al.
 (2020)\@BBCP.
-The vehicle routing is typically not considered in the electric bus scheduling, as buses usually have fixed stop points and timetables.
-After determining a finite set of candidate locations for potential EV charger deployment based on stop frequency, the number and types of chargers can be optimized according to trip patterns.
+After determining a finite set of locations for EV charger deployment, the numbers and types of chargers can be optimized based on committed trips and available buses.
+In prior studies, including \@BBOPcite\@BAP\@BBNHu et al.
+(2022)\@BBCP, \@BBOPcite\@BAP\@BBNLegault et al.
+(2025)\@BBCP, and \@BBOPcite\@BAP\@BBNGkiotsalitis et al.
+(2025)\@BBCP, the charging planning objective includes charging infrastructure costs, BEV purchase costs, and operational costs such as scheduling penalties and charging prices.
 \@BBOPcite\@BAP\@BBNHu et al.
-(2022)\@BBCP models a joint optimization problem for locating fast chargers at selected bus stops and optimizing schedules for opportunity charging.
-The model assumes that buses will be charged during layovers, and the cost takes into account the time-varying electricity price and penalty for additional waiting times.
+(2022)\@BBCP model a joint optimization problem for locating fast chargers at selected bus stops and optimizing schedules during bus layovers, and the objective function incorporates the time-varying electricity price and penalty for the additional waiting times.
 \@BBOPcite\@BAP\@BBNLegault et al.
-(2025)\@BBCP formulates a multi-period, integrated planning and scheduling problem.
-The paper models the electric bus timetable as a graph representing their charging and driving pattern, which is considered a deterministic, minimum-cost problem.
+(2025)\@BBCP formulates a multi-period integrated charging infrastructure planning and scheduling problem for an electric bus fleet.
+The model accounts for BEV and charger acquisition costs over multiple years, as well as the phased retirement of diesel buses.
 \@BBOPcite\@BAP\@BBNGkiotsalitis et al.
-(2025)\@BBCP optimize the EV charger location and number by allocating bus trips to their nearest chargers with available time slots, assuming that buses will be driven to the optimal charger location after the trip ends.
-1.2.2 EV charging scheduling with waiting time and abandonment
-The scheduling problem is a fundamental and widely studied combinatorial optimization problem, with extensive applications in areas such as supply chain management, service management, and healthcare \@BBOPcite\@BAP\@BBNFonseca et al.
-(2024), Delarue et al.
+(2025)\@BBCP optimize the mixed-type BEV chargers’ locations and numbers by allocating bus trips to their nearest chargers when they are available, assuming that buses are charged at the depots or the end of their trips.
+1.2.2 BEV Charging Scheduling
+The classical multi-depot vehicle scheduling problem takes a set of timetabled trips as input and requires that they all be performed by a single vehicle.
+The BEV charging scheduling problem is considered as a variant of this problem in which battery SoC constraints are imposed.
+The problem can be modeled as a network with nodes corresponding to discretized battery SoC levels and solved as a shortest path problem \@BBOPcite\@BAP\@BBNDe Vos et al.
 (2024)\@BBCP.
-EV charging can be viewed as a service that relies on a reusable resource—namely, the charging infrastructure.
-The EV charging scheduling problem could thus consider different attributes, such as charger availability, charging time, and waiting times for charging services.
-For instance, the delay or waiting time can be modeled as a monetary penalty in the objective function \@BBOPcite\@BAP\@BBNHu et al.
+The charging procedure can be viewed as a service scheduling if a fleet of vehicles is involved.
+The charging problem thus requires consideration of charger availability and waiting times (or queues) for charging services.
+For instance, the waiting time can be modeled as a monetary penalty in the objective function \@BBOPcite\@BAP\@BBNHu et al.
 (2022), Tao et al.
@@
 (2025)\@BBCP.
-When the waiting time exceeds a threshold, the vehicle that committed a time-critical task or trip might abandon charging and leave to find another charging spot, which can be represented as a scheduling problem with abandonment \@BBOPcite\@BAP\@BBNG.
-Chen et al.
-(2023)\@BBCP.
-Another critical aspect of EV charging is range anxiety, referring to drivers’ concern about depleting their battery before reaching a charging station or their destination.
-Empirical studies suggest that range anxiety may incentivize the installation of additional charging stations or the adoption of larger battery capacities \@BBOPcite\@BAP\@BBNLim et al.
+When the waiting time exceeds a threshold, the vehicle that initiated a time-critical task or trip may abandon charging and seek another charging spot, which can be modeled as abandonment.
+A unique aspect of BEV charging is range anxiety, the concern among drivers that their battery will deplete before reaching a charging station or their destination.
+Empirical studies show that range anxiety may incentivize the installation of additional charging stations or the adoption of larger battery capacities \@BBOPcite\@BAP\@BBNLim et al.
 (2015)\@BBCP.
@@
 \@BBOPcite\@BAP\@BBNTao et al.
-(2024)\@BBCP formulate the accumulated range anxiety level as a function of SoC depletion along the route and integrates the range anxiety into its objective function.
-1.2.3 Decision-dependent robust optimization
-In integrated charging infrastructure planning and scheduling problems, several uncertainties arise, including EV arrival and departure times, charging durations, and energy consumption.
-They can be modeled using stochastic optimization (SO), robust optimization (RO), and distributionally robust optimization (DRO).
-Among these formulations, RO typically ensures the highest level of reliability, making it particularly suitable for critical infrastructure planning.
-Different shapes of uncertainty sets, such as box or ellipse, lead to the robust counterparts and tractability \@BBOPcite\@BAP\@BBNBertsimas & Hertog (2022)\@BBCP.
+(2024)\@BBCP formulate the accumulated range anxiety level as a function of SoC depletion along the route and integrates this term into the objective function.
+1.2.3 Charging Event Characteristics and Uncertainty
+Charging events are usually characterized by charging duration and charging power (or energy), and are correlated with diverse features including charging type, hour of day, day of week, location, and vehicle type \@BBOPcite\@BAP\@BBNCui et al.
+(2023), Zhang et al.
+(2025)\@BBCP.
+\@BBOPcite\@BAP\@BBNCui et al.
+(2023)\@BBCP processed and analyzed more than 220,000 real-world charging records in Beijing.
+The authors clustered these charging events based on diverse features and structured them as a behavioral tree of fast charging.
+These clusters are labeled as low-, medium-, and high-priority based on dwell duration.
+In BEV charging scheduling problems, charging and dwell durations are typically unknown at the time of scheduling.
+These uncertainties can be modeled using stochastic optimization (SO), robust optimization (RO), and distributionally robust optimization (DRO).
+Among these formulations, RO typically provides the highest level of reliability, making it particularly suitable for critical infrastructure planning.
+Different shapes of uncertainty sets, such as boxes or ellipsoids, lead to different robust counterparts and affect tractability \@BBOPcite\@BAP\@BBNBertsimas & Hertog (2022)\@BBCP.
 \@BBOPcite\@BAP\@BBNHu et al.
-(2022)\@BBCP model the uncertainty of passengers’ boarding and alighting time during bus trips using box and budget uncertainty sets based on the RO formulation.
+(2022)\@BBCP model uncertainty in passengers’ boarding and alighting times during bus trips using box and budget uncertainty sets in an RO formulation.
 \@BBOPcite\@BAP\@BBNL.
 Chen et al.
-(2024)\@BBCP conduct EV charging scheduling considering the stochastic arrival time of different user groups, where an exponential cone programming formulation is introduced to approximate the upper bound of the stochastic problem.
+(2024)\@BBCP consider EV charging scheduling with stochastic arrival times of different EV user groups and introduce an exponential-cone programming formulation to approximate an upper bound.
 \@BBOPcite\@BAP\@BBNY.
-Chen & Liu (2023)\@BBCP developed a two-stage stochastic model for planning and operating shared autonomous EV systems, where scenarios represent uncertain user trip requests.
-An accelerated two-phase Benders decomposition-based algorithm is proposed for solving this two-stage programming problem.
-However, these constructed uncertainty sets model uncertainties independent of exogenous factors such as charging decisions.
-A number of charging installation trials have already shown that charging decisions are very likely to impact parking and waiting durations.
-For instance, parking duration is strongly correlated with the charging types (e.g., whether it is slow or fast charging) \@BBOPcite\@BAP\@BBNCui et al.
+Chen & Liu (2023)\@BBCP develop a two-stage stochastic model for planning and operating shared autonomous EV systems and create scenarios for uncertain user trip requests; an accelerated two-phase Benders decomposition-based algorithm is proposed to solve the resulting problem.
+\@BBOPcite\@BAP\@BBNWang et al.
+(2023)\@BBCP develop an ML-cluster-based, feature-driven ambiguity set for surgery scheduling using a DRO formulation; the clustering is based on classification and regression trees.
+1.2.4 Decision-dependent robust optimization for charging decisions
+With electrification, parking durations that previously involved no charging activities are likely to change after charger deployment.
+Large-scale charging-installation trials conducted in multiple countries have demonstrated that charging decisions significantly affect BEV parking and waiting durations \@BBOPcite\@BAP\@BBNCui et al.
 (2023), Zhang et al.
 (2025)\@BBCP.
-This could be modeled as the decision-dependent uncertainty set, and its uncertainty is endogenous \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP.
-\@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP demonstrates that, in general, RO problems with decision-dependent uncertainty are NP-complete and require either global optimization or reformulation techniques such as the big-M method.
-Similarly, \@BBOPcite\@BAP\@BBNNoyan et al.
-(2022)\@BBCP introduces a decision-dependent ambiguity set in a machine scheduling problem and examines how different types of uncertainty dependence affect the model formulation.
-Their findings show that incorporating affine dependency on decision variables can reduce nonlinear constraints through the use of McCormick envelopes and the big-M approach.
-Decision Variables User Behavior Uncertainty formulation
+Parking (or charging) duration can thus be modeled using a decision-dependent uncertainty set that depends on endogenous factors, such as charging decisions.
+\@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP demonstrate that, for simple robust linear problems, RO-DDU is NP-complete and requires either global optimization or reformulation techniques such as the big-M method.
+\@BBOPcite\@BAP\@BBNLuo & Mehrotra (2020)\@BBCP study decision-dependent DRO models with five types of ambiguity sets, where an instantaneous decision directly impacts right-hand-side uncertainty, including moments and the radius of ambiguity sets.
+\@BBOPcite\@BAP\@BBNNoyan et al.
+(2022)\@BBCP propose a decision-dependent ambiguity set for a machine scheduling problem and show that introducing affine dependence on decision variables can substantially reduce the number of constraints induced by McCormick envelopes and big-M reformulations.
+However, existing studies are largely limited to settings with instantaneous decisions, whereas in practice, uncertainty is influenced by aggregated and expected decision outcomes.
+Decision Variables User Behaviors Uncertainty formulation
 Location Number Scheduling Type Waiting time Range anxiety
@@
 \@BBOPcite\@BAP\@BBNGkiotsalitis et al. (2025)\@BBCP D
-Our paper RO
-  • Notes: D = Deterministic model; RO = Robust Optimization; SO = Stochastic Optimization.
-All the aforementioned work about charging infrastructure planning and scheduling is presented in Table 1.2.3. Compared to these works, our paper makes the following contributions:
-  • We developed a two-stage EV charger planning model that optimizes the charger location, quantity, type, and scheduling.
-The model introduces a scheduling-with-abandonment formulation that reflects the industrial site’s strict operational rules and captures the impacts of overnight parking, range anxiety, and waiting times on the charging scheduling.
-  • We developed a decision-dependent uncertainty set to capture the variability in parking durations and their differences across charger types (i.e., slow and fast chargers).
-Although decision-dependent robust optimization models are typically nonlinear, we demonstrated that the proposed affine dependence on the binary charging decision can be linearized, allowing the problem to be formulated as a MILP.
-  • Considering the high computation cost of the scheduling problem over a long optimization horizon, we designed a ‘fix-and-optimize’ heuristic method to forecast parking intervals with high charging demands in the scheduling.
-Using the representative days or iterative planning by month, the model achieves an optimality gap of less than 0.1% within a reasonable computation time under four uncertainty scenarios.
+Our paper RO-DDU
+  • Notes: D = Deterministic model; RO = Robust Optimization; SO = Stochastic Optimization; RO-DDU = Robust Optimization with Decision-Dependent Uncertainty.
+Table 1: Literature Review Summary
+All the aforementioned work on charging infrastructure planning and scheduling is summarized in Table 1. Compared with these works, our paper makes the following contributions:
+  • We develop an integrated charger planning and scheduling model that optimizes charger locations, quantities, and types, and jointly determines schedules given the selected infrastructure.
+The model introduces a scheduling-with-abandonment formulation that reflects the industrial site’s strict operational rules and captures the impacts of overnight parking, range anxiety, and waiting times.
+  • We develop a feature-based, decision-dependent uncertainty set to capture the variability of charging durations and their dependence on charging decisions.
+Leveraging ML techniques and contextual information, we restructure time–zone features to preserve the calendar effects of truck trips and group parking durations to efficiently set the bounds of the uncertainty sets.
+  • Given the limited discussion in the literature on dependency assumptions in decision-dependent uncertainty problems, we distinguish between two cases in which charging duration depends on either instantaneous charging decisions or cluster-level expected charging decisions.
+We derive tractable reformulations for both cases and discuss the resulting outcomes and managerial implications.
+  • Considering the high computational cost of solving the scheduling problem over a long optimization horizon, we design a “fix-and-optimize” heuristic to forecast parking intervals with potentially high charging demand.
+Using representative days or a rolling horizon, the model achieves an optimality gap of less than 0.1% within a reasonable computation time under different scenarios.
 2 Problem formulation
 2.1 Problem description
-The problem is formulated as a two-stage model: the first stage determines charger installations, and the second stage schedules truck charging based on the installed infrastructure.
-We consider a fleet of electric light-duty trucks $i\in\mathcal{I}$ operating around an industrial site based on the fixed stop points and schedules.
-As shown in Figure 1, a truck $i$ can be charged during its available half-hour parking time slots $\mathcal{S}_{i,d,t}$ at time $t$ on day $d$ at a charging zone $z$.
-One truck cannot appear in more than one zone at a time.
-Figure 1: The studied multi-zone charging scheduling problem for the light-duty truck fleet; At any half-hour parking time slot $\mathcal{S}$ shown in light green, a truck can either choose to wait, abandon charging, or charge.
-The effective parking durations pppp are shown in dark green.
-Within a parking slot ${S}_{i,d,t}$, a truck may either wait to be charged, abandon charging, or charge directly using one of the available charger types $j\in\mathcal{J}$.
-In each parking time slot, the effective parking duration is less than half an hour, which restricts the maximum amount of energy that can be charged for this truck during this parking time window.
-Across the optimization horizon, the total energy of each truck scheduled to be charged must satisfy the truck’s trip-related energy consumption, considering the battery energy balance, battery SoC constraints, and charging power constraints.
-The traveling distance and energy consumption during parking slots ${S}_{i,d,t}$ is not considered.
-Furthermore, at any time $t$ on the day $d$, the number of chargers simultaneously used in zone $z$ in type $j$ cannot exceed the number of this type of chargers installed in the zone, $x_{z}^{j}$.
-2.2 Set, index, and parameters
+In the joint planning and scheduling problem, we consider a fleet of electric light-duty trucks $i\in\mathcal{I}$ operating around an industrial site based on the fixed stop points and schedules during the time period $\mathcal{T}$.
+Figure 1: The studied multi-zone charging scheduling problem for the light-duty truck fleet.
+At any half-hour parking time slot $\mathcal{S}$ shown in light green, A truck can choose to wait, abandon charging, or charge.
+The effective parking durations $pp$ are shown in dark green.
+As shown in Figure 1, a truck $i$ can either wait to be charged, abandon charging, or charge directly using one of the available charger types $j\in\mathcal{J}$ during its available half-hour parking time slots $\mathcal{S}_{i,d,t}$ at time $t$ on day $d$ in charging zone $z$.
+A truck cannot appear in more than one zone at a time.
+In each parking time slot, the effective parking duration is less than half an hour, which restricts the maximum amount of energy that can be charged during this window.
+Across the optimization horizon, the total energy scheduled for each truck must satisfy the truck’s trip-related energy consumption, subject to the battery energy balance, battery SoC constraints, and charging power constraints.
+Traveling distance and energy consumption during parking slots $\mathcal{S}_{i,d,t}$ are not considered.
+Furthermore, at any time $t$ on day $d$, the number of chargers of type $j$ used simultaneously in zone $z$ cannot exceed the number of installed chargers of that type in the zone, $x_{z}^{j}$.
+2.2 Set, index and parameters
   • $\mathcal{J},j$: Set, index of EV charger types.
@@
   • $\sigma$: Standard deviation of parking durations by truck $i$ at hour $t$ on day $d$ at zone $z$ across all the optimized days.
-  • $\Delta t$: Duration (mins) in each parking time interval
+  • $\Delta t$: Duration (mins) in each parking time interval.
 2.3 Decision variables
@@
   • $pp_{i,j,z,t}\in\mathbb{R}^{+}$: Parking duration for truck $i$ at time $t$ in zone $z$ using charging type $j$.
+  • $\bar{Y}_{c}^{j}\in\mathbb{R}^{+}$: The expected cluster-wise charging decision with the cluster $c$.
 2.5 Objective function
-The objective function consists of two stages: the charging planning and scheduling stages.
-The planning stage considers multiple types of charging infrastructure, such as slow and fast chargers.
-In addition to power capacity, we also model the SoC charging range for each type of charger, which is detailed in the following section.
-The operational part includes penalties for low battery SoC, which induces range anxiety and affects charging time during scheduling.
+The objective function includes the charging planning and scheduling.
+The charger planning accounts for multiple charging infrastructure types, including slow and fast chargers.
+In addition to power capacity, we also model the SoC charging range for each type of chargers, which are detailed in the following section.
+The scheduling part includes penalties for low battery SoC, inducing the range anxiety and the total charging time.
 […]
@@
 This variable is greater than zero (i.e., $v_{i,d,t}>0$) if and only if the battery SoC level falls below 30 %.
-The charging time penalty is calculated as the sum of all the charging times across all zones and the time horizon to prevent low charger utilization.
+The charging-time penalty is calculated as the sum of charging times across all zones and the time horizon to prevent low charger utilization (i.e., low average charging power).
 2.6 Constraints for charging planning and scheduling
@@
 Here, $M_{1}$ is a large constant, and the effective range of $M_{1}$ is presented in Appendix 6.
-The parameter $\epsilon$ is a small positive number, and it ensures that the indicator equals one only when the battery SoC is strictly greater than 30%, thereby guaranteeing a unique waiting time when the battery SoC is exactly 30%.
+The parameter $\epsilon$ is a small positive number that ensures the indicator equals 1 only when the battery SoC is strictly greater than 30%, thereby guaranteeing a unique waiting time when the battery SoC is exactly 30%.
 2.6.2 Abandonment and waiting time constraints
@@
 $M_{2}$ is a large constant to relax the waiting time constraint when $a_{i,d,t}$ equals to 1, in other words, when the trucks abandon charging.
-Notably, the maximum waiting time is a variable and it depends on the truck SoC due to the range anxiety and whether it is parked in special zones that allow for the overnight charging.
-Based on experiences from the industrial partner, we formulate the following waiting and abandonment rules, considering range anxiety and overnight charging in the special zones.
+Notably, the maximum waiting time varies with the truck’s range anxiety and with whether it is parked in special zones that permit overnight charging.
+Based on the industrial partner’s experience, we formulate the following waiting and abandonment rules, accounting for range anxiety and overnight charging in special zones as follows.
 As constraint (11), if the SoC is greater than or equal to 30%, no waiting time is allowed as the driver has no range anxiety.
-If the SoC of a truck is below 30%, this will result in the driver’s range anxiety and would extend their waiting time to 0.5 hours before abandoning charging.
+If the SoC of a truck is below 30%, this will result in the driver’s range anxiety and would extend their waiting time to half an hour before abandoning charging.
 Let the set $S^{Z}$ denote the parking time slots at any of the special zones for overnight charging.
@@
 Constraint (16) sets the maximum charging power limit considering the charging type $P^{j}$ and corresponding charging efficiency $\eta^{j}$.
-Constraint (17) regulates the charging time limit in each time window, which is greater than 5 mins and less than 30 minutes.
+Constraint (17) regulates the charging time limit in each time window, which is greater than 5 mins and less than 30 mins.
 Finally, constraint (18) restricts battery SoC within the maximum and minimum values.
@@
 The pair $(d^{\prime},t^{\prime})$ represents the next time step after $(d,t)$, defined as $(d,t+1)$ if t<T1t<T-1, or $(d+1,0)$ if t=T1t=T-1 and d<D1d<D-1, where $T$ is the number of time steps per day and $D$ is the number of days.
-2.6.6 Decision-dependent uncertainty sets of parking duration
-The effective charging duration $\hat{pp}$ is stochastic and influenced by the charging decision $y_{i,d,t}^{j}$, which is represented as binary variables.
-We constructed a box-shaped, decision-dependent uncertainty set for truck $i$ in zone $z$ at hour $t$.
-This is formulated as a robust linear constraint with affine decision dependence in terms of charging decisions, $y_{i,t,z}^{j}$.
-$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}\cdot P^{j}\cdot\Delta t\cdot y_{i,d,t}^{j}\cdot\hat{pp}_{i,t,z},\quad\hat{pp}\in\Phi(y_{i,t,z}^{j})\quad\forall i,d,t\in\mathcal{S}$ (21)
-The box-shaped uncertainty set $\Phi(y_{i,t,z}^{j})$ is defined as,
-$\Phi=\left\{\hat{pp}\in\mathbb{R}^{i\times t}\ \Bigg|\ \frac{5}{30}\leq\hat{pp}\leq 1,\ \mu_{i,t,z}(y_{i,t,z}^{j})-\sigma_{i,t,z}(y_{i,t,z}^{j})\leq\hat{pp}\leq\mu_{i,t,z}(y_{i,t,z}^{j})+\sigma_{i,t,z}(y_{i,t,z}^{j})\right\}$ (22)
-Where the bounds of the uncertainty set depend on the the minimum and maximum charging time, and calculated mean $\mu_{i,t,z}$ and standard deviation $\sigma_{i,t,z}$ associated with the charging decisions.
-The nominal moments are extracted from the historical data.
-Many charging trial data shows that fast charging could reduce the uncertainty of parking duration as the charging time becomes short.
-Therefore, we introduced a factor $\gamma$ to represent the reduction in uncertainty resulting from the fast charger usage.
-$\mu_{i,t,z}(y_{i,t,z}^{j}):=\hat{\mu}_{i,t,z}\left(y_{i,t,z}^{0}+y_{i,t,z}^{1}\right)$ (23)
-$\sigma_{i,t,z}(y_{i,t,z}^{j}):=\hat{\sigma}_{i,t,z}\left(\gamma^{0}\cdot y_{i,t,z}^{0}+\gamma^{1}\cdot y_{i,t,z}^{1}\right)$ (24)
-Where $y^{0}$ and $y^{1}$ indicate charging decisions using slow or fast chargers, respectively.
-Notably, the decision-dependent RO is NP-complete, non-linear due to the multiplication of variables \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP.
-Here we proved that the proposed formulation is a special case that can be formulated as a robust linear constraint (33), and therefore the model can be formulated as a MILP problem.
-The proof is presented in Appendix 8.
-$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}\cdot P^{j}\cdot\Delta t\cdot(\hat{\mu}_{i,t,z}\cdot y_{i,d,t}^{j}-\hat{\sigma}_{i,t,z}\cdot\gamma^{j}\cdot y_{i,t,z}^{j}),\quad\forall i,d,t\in\mathcal{S}$ (25)
 2.6.6 Total number of chargers in each type and zone
@@
 Where $\mathbbm{1}$ is a binary indicator to represent whether truck $i$ is at zone $z$ at time $t$ of day $d$.
-3 Robust planning using the whole year dataset
-The planned EV infrastructure must remain robust across the year-round fleet operations.
-We compare two approaches to solving based on the whole-year dataset: day selection (DS) and iterative planning and scheduling (IPS).
-  • DS: We selected a certain number of representative days based on statistical criteria, such as the total fleet distance, the number of moving trucks, the total fleet energy consumption, and the total stopping time during a day.
-We also included days around the representative days to preserve the trip and battery cycle continuity.
-  • IPS: We first decomposed the full-year dataset into monthly subsets, allowing several overlapping days between months.
-The algorithm then plans charger deployment for the entire year by iteratively integrating scheduling and monthly planning.
-It starts with an initial charger allocation derived from the dataset and processes each month sequentially.
-If the existing infrastructure can satisfy the charging demand of this month, only charging scheduling is performed.
-Otherwise, the algorithm jointly determines the charging schedule and the additional chargers required, updates the infrastructure, and carries the new configuration forward to subsequent months.
-4 Fix-and-optimize heuristic based on high charging demand forecasts
-Directly solving the proposed planning and scheduling model can be computationally demanding with a long optimization horizon.
-In this paper, we propose a fix-and-optimize heuristic designated for the scheduling problem based on high charging demand forecasts.
-The fix-and-optimize heuristic has been applied to timetable \@BBOPcite\@BAP\@BBNFonseca et al.
+2.7 Decision-dependent robust formulation considering parking duration uncertainty
+The stochastic effective parking duration, denoted by the random variable $\hat{pp}_{i,d,t}$, is unknown at the time of scheduling.
+We assume that this duration is endogenous: it depends on the charging decision $y_{i,d,t}^{j}$.
+Specifically, we model the realized parking duration $(\hat{pp})_{i,d,t}^{\prime}$ as an affine transformation of the intrinsic (base) random duration $\hat{pp}_{i,d,t}$.
+Let $k^{j}\geq 1$ be a scaling constant associated with charger type $j$ (e.g., representing the reduction in charging time due to fast charging). The decision-dependent duration is defined as:
+$(\hat{pp})_{i,d,t}^{\prime}:=\hat{pp}_{i,d,t}\cdot\left(1-\sum_{j\in\mathcal{J}}\frac{y_{i,d,t}^{j}}{k^{j}}\right)$ (22)
+Since we enforce single charger usage ($\sum y^{j}\leq 1$), the term in the parentheses acts as a scalar scaling factor.
+2.7.1 Feature-driven cluster-wise uncertainty sets
+Assuming that we have a total number of $N$ historical records of parking duration, each is indexed by $\omega\in\Omega$, $\Omega=\{1,2,...,N\}$, and records a truck’s parking duration $pp^{\omega}$ and its feature vector $v^{\omega}\in\mathcal{V}$, where $\mathcal{V}$ is the feature space.
+Each feature vector encodes information for each parking duration, including contextual and numerical variables such as half-hour of day, day of year, and zone.
+The fleet information describes the trucks’ operational purposes and is categorized into eight types, such as drill-and-blast or production activities.
+For instance, the feature vector of a parking duration $\omega$ can be represented as,
+$\boldsymbol{v}^{\omega}=$ $\{1:(00:00),20:(\text{The 20th day of the year 2023}),$
+$4:(\text{Zone 4}),3:(\text{Production team}),255:(\text{Truck ID 255})\}$
+We found that parking duration is strongly correlated with fleet groups, half-hour intervals, and zones, but only slightly related to the day of the year.
+This motivates us to partition the parking duration into $c\in\mathbb{N}$ clusters to represent its stochasticity.
+Given the calendar effect of parking events, we propose using hierarchical clustering by reordering the feature vectors for each truck.
+We compute the average parking duration for each half-hour zone combination, aggregated across all days of the year.
+These averaged parking durations under half-hour zone combinations serve as the feature vectors for individual trucks.
+Let the newly arranged feature vector be $(\boldsymbol{v}^{i})^{{}^{\prime}}$.
+Using these time–zone features alongside fleet information, we perform hierarchical clustering to group trucks, using Euclidean distance as the similarity measure.
+$(\boldsymbol{v}^{i})^{{}^{\prime}}=$ $\{1:(00:00;\text{Zone 1}),2:(00:30;\text{Zone 1}),...,N_{t}\cdot N_{z}:(24:00;\text{Zone 8})\}$
+We next model the distribution of uncertain parking duration for all trucks or truck groups and construct the corresponding feature-driven uncertainty set.
+Assuming that the parking duration of each cluster is identically and independently distributed, we could calculate the empirical mean $\mu_{c}$ and variance $\sigma_{c}^{2}=\frac{\sum_{\omega\in\Omega_{c}}(\hat{pp}^{\omega}-\mu_{c})^{2}}{N_{c}}$.
+Considering the inherent bounds of parking duration, we build a box-shaped uncertainty set as follows,
+$\mathbb{P}:=\Bigg\{\hat{pp}\;\Big|\;\begin{array}[]{l}\mathbb{E}_{\mathbb{P}}[\hat{pp}]=\boldsymbol{\mu}_{pp},\\[2.0pt] \mathbb{E}_{\mathbb{P}}\!\left[(\hat{pp}-\boldsymbol{\mu}_{pp})(\hat{pp}-\boldsymbol{\mu}_{pp})^{\mathsf{T}}\right]=\boldsymbol{\Sigma}_{pp},\\[2.0pt] \lVert\hat{pp}\rVert_{\infty}\leq\Delta_{c}\end{array}\Bigg\}$ (23)
+where $\boldsymbol{\mu}_{pp}=(\mu_{c}(\boldsymbol{v}))_{c\in\mathcal{C}}$, and the covariance matrix $\boldsymbol{\Sigma}_{pp}=\text{diag}(\sigma_{c}^{2}(\boldsymbol{v})_{c\in\mathcal{C}})$ are feature-driven.
+The values of $\Delta_{c}$ are based on $\boldsymbol{\mu}_{pp}$ and $\boldsymbol{\Sigma}_{pp}$, which control the conservativeness of our formulation.
+2.7.2 Decision-dependent uncertainty sets and moment definitions
+We could construct a decision-dependent uncertainty set $\mathbb{P}$ (23) based on the first and second moments.
+Let $\boldsymbol{\mu}_{c}$ and $\boldsymbol{\Sigma}_{c}$ denote the empirical mean vector and covariance matrix of the base parking duration for a cluster $c$, derived from historical data without charging interventions.
+We set $\Delta_{c}$ to be $\boldsymbol{\mu}_{c}\pm\boldsymbol{\sigma}_{c}$.
+[…]
+Example 1: Affine dependency on instantaneous decisions.
+The moments are scaled directly by the specific decision for that parking slot.
+As $y_{i,d,t}^{j}$ is a “here-and-now” decision variable fixed prior to uncertainty realization, the adjusted moments are:
+$\mu_{c}(y_{i,d,t})$ $=\hat{\mu}_{c}\left(1-\sum_{j\in\mathcal{J}}\frac{y_{i,d,t}^{j}}{k^{j}}\right)$ (24)
+$\Sigma_{c}(y_{i,d,t})$ $=\hat{\Sigma}_{c}\left(1-\sum_{j\in\mathcal{J}}\frac{y_{i,d,t}^{j}}{k^{j}}\right)^{2}$ (25)
+Example 2: Affine dependency on cluster-level expected decisions.
+Here, the scaling factor depends on the aggregate behavior of the fleet within a cluster $c$.
+Let $N_{c}$ be the total number of parking slots in cluster $c$.
+We define the empirical expected value of charging decision $\bar{Y}_{c}$ as:
+$\bar{Y}_{c}^{j}:=\frac{1}{N_{c}}\sum_{(i,d,t)\in c}y_{i,d,t}^{j}$ (26)
+Note that $\bar{Y}_{c}$ is a linear function of decision variables, not a stochastic expectation.
+Proposition 1 Given that the empirical expected value $\bar{Y}_{c}$ for a cluster of parking duration $c$, the moments for any instance in this cluster are scaled uniformly:
+$\mu_{c}(\bar{Y}_{c}^{j})$ $=\hat{\mu}_{c}\left(1-\sum_{j\in\mathcal{J}}\frac{\bar{Y}_{c}^{j}}{k^{j}}\right)$ (27)
+$\Sigma_{c}(\bar{Y}_{c}^{j})$ $=\hat{\Sigma}_{c}\left(1-\sum_{j\in\mathcal{J}}\frac{\bar{Y}_{c}^{j}}{k^{j}}\right)^{2}$ (28)
+Proof of Proposition 1 Please see 7.1 of the Supporting information.
+2.7.3 RO-DDU constraint and reformulations
+To ensure schedule feasibility, we enforce a RO-DDU constraint on the charging power limit:
+[…]
+[…]
+Theorem 1.
+Example 1: Affine dependency on instantaneous decisions.
+Constraint (30) is the reformulation of the linear power limit constraint (29) under the robust uncertainty set with affine dependency on instantaneous decisions $\boldsymbol{y}^{j}$.
+This formulation preserves the linearity of binary charging decisions, keeping the problem tractable.
+$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}P^{j}\Delta t\left[\boldsymbol{\hat{\mu}}_{c}^{\intercal}\left(1-\frac{1}{k^{j}}\right)\boldsymbol{y}^{j}-\left(1-\frac{1}{k^{j}}\right)\boldsymbol{\hat{\Sigma}}^{1/2}_{c}\boldsymbol{y}^{j}\right]$ (30)
+Example 2: Affine dependency on cluster-level expected decisions.
+After applying decision-dependent moments (27) and (28), constraint (31) is the reformulation of the linear power limit constraint (29) under the robust uncertainty set with affine dependency on cluster-level expected decisions $\bar{Y}_{c}$.
+$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}P^{j}\Delta t\left[\boldsymbol{\hat{\mu}}_{c}^{\intercal}\left(1-\frac{\bar{Y}_{c}^{j}}{k^{j}}\right)\boldsymbol{y}^{j}-\left(1-\frac{\bar{Y}_{c}^{j}}{k^{j}}\right)\boldsymbol{\hat{\Sigma}}^{1/2}_{c}\boldsymbol{y}^{j}\right]$ (31)
+This formulation is nonlinear because it includes the product of the binary variable $\boldsymbol{y}^{j}$ and the continuous variable $\bar{Y}_{c}$.
+Proof of Theorem 1. Please see 7.2 of the Supporting information.
+2.8 Cluster-based Linearization (Example 2)
+While statistically rigorous, constraint (31) involves the product of the continuous variable $\bar{Y}_{c}$ (the sum of binary variables normalized by $N_{c}$) and the binary decision variable $y^{j}$, introducing bilinear terms.
+To solve this within a MILP framework, we introduce an auxiliary variable $Z_{i,d,t}^{j}=\bar{Y}_{c}^{j}\cdot y_{i,d,t}^{j}$.
+Since $\bar{Y}_{c}\in[0,1]$ and $y_{i,d,t}^{j}\in\{0,1\}$, this product can be exactly linearized using the McCormick envelopes \@BBOPcite\@BAP\@BBNMcCormick (1976)\@BBCP:
+$Z_{i,d,t}^{j}\leq\bar{Y}_{c}^{j}$ (32)
+$Z_{i,d,t}^{j}\leq y_{i,d,t}^{j}$ (33)
+$Z_{i,d,t}^{j}\geq\bar{Y}_{c}^{j}-(1-y_{i,d,t}^{j})$ (34)
+$Z_{i,d,t}^{j}\geq 0$ (35)
+Substituting $\boldsymbol{Z}^{j}$ into (31) renders the constraint linear in the optimization variables, albeit at the cost of increasing the problem size. We obtain the constraint (36).
+$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}P^{j}\Delta t\left[\boldsymbol{\hat{\mu}}_{c}^{\intercal}\left(1-\frac{1}{k^{j}}\right)\boldsymbol{Z}^{j}-\left(1-\frac{1}{k^{j}}\right)\boldsymbol{\hat{\Sigma}}^{1/2}_{c}\boldsymbol{Z}^{j}\right]$ (36)
+3 Solution Algorithm
+3.1 Robust planning using the whole year dataset
+The planned EV infrastructure must remain robust across year-round fleet operations, that is, capable of meeting trip-related energy consumption, given the limited parking duration.
+We compare two planning results under different scheduling operations, based on the selective or full dataset:
+In terms of data usage, data can be categorized as day selection (DS) or full year (FY).
+For DS, we selected a subset of representative days based on statistical criteria, including total fleet distance, number of moving trucks, total fleet energy consumption, and total stopping time per day.
+We also include days around the representative days to preserve the trip and battery-cycle continuity.
+In contrast, the FY method uses the full-year dataset directly.
+For scheduling operations, we consider the hindsight (HS) and iterative planning and scheduling (IPS) approaches.
+The HS method assumes perfect foresight and performs one-shot scheduling.
+The IPS method is similar to rolling-horizon scheduling.
+The full-period dataset is first decomposed into monthly subsets, allowing several overlapping days between periods.
+The algorithm then plans charger deployment for the entire period by iteratively integrating scheduling and planning at the sub-period level.
+It starts with an initial charger allocation derived from the dataset and processes each sub-period sequentially.
+If the existing infrastructure can satisfy the charging demand of this sub-period, only charging scheduling is performed.
+Otherwise, the algorithm jointly determines the charging schedule and the additional chargers required, updates the infrastructure, and carries the new configuration forward to subsequent sub-periods.
+Figure 4: The fix-and-optimize heuristic where fast-charging scheduling is assigned and fixed based on the forecast high charging demands
+The proposed planning and scheduling model is difficult to solve when the optimization horizon is long, as in hindsight scheduling.
+We propose a fix-and-optimize heuristic for the scheduling problem based on high-charging-demand forecasts.
+It has been applied to the timetable \@BBOPcite\@BAP\@BBNFonseca et al.
 (2024)\@BBCP and job-shape scheduling \@BBOPcite\@BAP\@BBNLi et al.
-(2025)\@BBCP problems, where part of the solution is fixed while the remaining variables are free to form subproblems.
-In this paper, as presented in Fig.
-4, we identified the parking intervals with high charging demand at an early stage based on the samples from the uncertainty set and pre-assigned fast-charging schedules, thereby accelerating the overall optimization process.
-Figure 4: The fix-and-optimize heuristic where fast-charging scheduling is pre-assigned and fixed based on the forecast high charging demands
-As shown in Algorithm 1, we first initialize the master problem using the starting values $pp_{i,d,t}$ to solve the model.
-Next, we sample the stochastic parking duration $\hat{pp}_{i,d,t}$ from the uncertainty set $\Phi(\mu,\epsilon)$ to replace the initial values and compute the upper bounds of the charging power based on these samples.
-We then compare the required charging power against the updated upper bounds to identify any violations.
-A violation occurs if the required charging power exceeds the upper bound by more than a threshold $\kappa^{v}$ and the parking window does not belong to a special zone.
-The threshold $\kappa^{v}$ is initialized at two and decays exponentially across iterations.
-When a violation is detected, a fast-charging scheduling constraint $y_{i,1,d,t}=1$ is added for that parking window.
-These constraints are subsequently incorporated into the master problem, after which the model is reset and resolved from the first step.
-Input: Parking duration samples $\hat{pp}_{i,d,t}$ drawn from the uncertainty set $\Phi(\mu,\epsilon)$
-Output: Charger installation decisions $x_{z}^{j}$
+(2025)\@BBCP, where part of the solution is fixed while the remaining variables are free to form subproblems.
+As shown in Fig.
+4, we first identified parking intervals with high charging demand and pre-assigned fast-charging schedules, thereby accelerating the overall optimization process.
+3.2 Fix-and-optimize heuristic based on high charging demand forecasts
+Input: Master problem $\mathcal{M}$ with deterministic parking duration $pp_{i,d,t}$; Parking duration clusters $c(i,t,z)\in\mathcal{C}$
+Output: BEV charger installation decisions $x_{z}^{j}$
 repeat
-   Solve the proposed optimization problem with deterministic $pp_{i,d,t}$ as the initial input (master problem); $\triangleright$ Warm start if the solution of the master problem available;
-   Obtain scheduling outcomes $y_{i,j,d,t}$ and charging profiles $p^{ch}_{i,j,d,t}$;
-  Sample parking durations $\hat{pp}_{i,d,t}$ from the uncertainty set $\Phi(\mu,\epsilon)$;
- for $i\in\mathcal{I},\;t\in\mathcal{T},\;d\in\mathcal{D},\;j\in\mathcal{J}$ do
-      Compute the new upper bound of charging power:
-$p^{ch^{\prime}}_{i,j,d,t}\leftarrow\sum_{j\in\mathcal{J}}\Delta t\cdot P^{j}\cdot y_{i,j,d,t}\cdot\hat{pp}_{i,d,t}\cdot\eta^{j}\;$
-    if $\big(p^{ch}_{i,0,d,t}\geq p^{ch^{\prime}}_{i,0,d,t}+\kappa^{v}\big)\land(\forall i,d,t\in\overline{\mathcal{S}^{Z}})$ or $p^{ch}_{i,1,d,t}\geq p^{ch^{\prime}}_{i,1,d,t}$ then
+   Solve the master problem $\mathcal{M}$; $\triangleright$ Warm start at the first step if the solution available;
+   Obtain scheduling outcomes $\mathbf{y}=(y^{j}_{i,d,t})_{(i,d,t)\in\mathcal{S},j\in\mathcal{J}}$ and charging power $\mathbf{p}^{ch}=(p^{ch})_{(i,d,t)\in\mathcal{S}}$;
+ for $(i,d,t)\in c\;,j\in\mathcal{J}$ do
+      Compute the lower bound of parking durations $\hat{pp}_{i,d,t}=\mu_{c}-\sigma_{c}$ for the cluster $c$;
+    $\triangleright$ Applying scaling factors $k^{j}$ if RO-DDU ;
+      Compute the updated upper bound of charging power $p^{ch^{\prime}}_{i,j,d,t}$
+$p^{ch^{\prime}}_{i,j,d,t}\leftarrow\sum_{j\in\mathcal{J}}\Delta t\cdot P^{j}\cdot y^{j}_{i,d,t}\cdot\hat{pp}_{i,d,t}\cdot\eta^{j}\;$
+    if Violations are detected then
          Impose the fast-charging constraint $y_{i,1,d,t}=1$ in the master problem;
 until The maximum iteration limit is reached;
-Post-process the solution by removing fast charging results in zones where the number of fast chargers exceeds two;
-Re-solve the optimization problem using the sampled values $\hat{pp}_{i,d,t}$ and incorporating the charger installation constraints derived from the processed solution;
-Algorithm 1 Fix-and-optimize heuristic under effective parking duration uncertainty
+Post-process the solution by discarding part of the solution and re-solve the problem with $\hat{pp}_{i,d,t}$;
+Algorithm 1 Fix-and-optimize heuristic under parking duration uncertainty
+As illustrated in Algorithm 1, we first initialize the master problem with the starting values $pp_{i,d,t}$.
+Next, we compute the lower bound of parking duration $\hat{pp}_{i,d,t}$ based on the cluster $c$ to replace the initial values and compute the updated bounds of the charging power.
+We then compare the updated upper bounds to the required charging power in the master problem $\mathcal{M}$ to identify any violations.
+A violation occurs if the required charging power exceeds the upper bound by more than a threshold $\kappa^{v}$ and the parking window does not belong to a special zone.
+The threshold $\kappa^{v}$ is initialized at 2 and decays exponentially across iterations.
+When a violation is detected, a fast-charging scheduling constraint $y_{i,1,d,t}=1$ is added for that parking window.
+These constraints are subsequently incorporated into the master problem, after which the model is resolved from the first step.
 This iterative process continues until the maximum number of iterations is reached.
-Afterward, we post-processed the optimization results by removing zones with more than two fast chargers.
-The remaining fast charger installations were treated as hard constraints, while the solutions for slow chargers were used as lower bounds.
+Afterward, we post-processed the optimization results by discarding part of the solution, for instance, a high number of fast chargers in one zone.
+The remaining solutions for fast and slow chargers are then used as lower bounds to resolve the model.
 Ultimately, the installation plan for all zones and charger types was determined.
 4 Application
-We applied our model to an open-pit mining site with 203 light-duty trucks in operation, which are planned to be replaced by battery EVs.
+We applied our model to an open-pit mining site with 203 light-duty trucks in operation, planned for replacement with BEVs.
 Seven special-use trucks, such as ambulances, that will not be electrified, along with another seven trucks dispatched fewer than 10 days, are excluded from the study.
 The remaining 189 trucks will be replaced with electric trucks equipped with 100 kWh batteries.
-We consider the following two types of chargers, as listed in Table 1.
-The capital cost of fast chargers includes the installation cost.
-We set the penalty of low SoC and charging time to be one in the objective function.
-The charging efficiency could vary with the charging type, outdoor temperature, and charging power \@BBOPcite\@BAP\@BBNSevdari et al.
+We consider the following two types of BEV chargers, as listed in Table 2.
+The capital cost of fast chargers does not include installation costs, as installation is location-dependent.
+We set the penalties for low SoC and charging time to be equal in the objective function.
+Although the charging efficiency could vary with the charging type, outdoor temperature, and charging power \@BBOPcite\@BAP\@BBNSevdari et al.
+(2023)\@BBCP, this paper uses a fixed value for charging efficiency.
+Overall, fast charging is more efficient than slow charging because it uses a high-efficiency rectifier located outside the truck \@BBOPcite\@BAP\@BBNArena et al.
 (2024)\@BBCP.
-This paper uses a fixed value for charging efficiency.
-Overall, fast charging tends to be more efficient than slow charging, as it uses a high-efficiency rectifier located outside the truck \@BBOPcite\@BAP\@BBNArena et al.
-(2024)\@BBCP.
-We therefore set the charging efficiency of slow and fast charging to be 90% and 95%.
+We therefore set the charging efficiencies for slow and fast charging to 90% and 95%.
 Specifications Slow charger Fast charger
-Capital costs (\$/charger) 1,500 \@BBOPcite\@BAP\@BBNFord (2025)\@BBCP 38,000 \@BBOPcite\@BAP\@BBNPrimecomtech (n.d.)\@BBCP
+Capital costs (\$/charger) 1,500 \@BBOPcite\@BAP\@BBNFord (2025)\@BBCP 38,000 \@BBOPcite\@BAP\@BBNPrimecomtech (2025)\@BBCP
 Maximum charging rates (kW) 11.2 150
@@
 The charging planning process begins by identifying potential charging sites based on the stop frequency of all trucks (Section 4.1).
-Next, the parking window, effective parking duration, and energy consumption for each truck are extracted (Section 5.2).
-Using this information, the DS and IPS methods are applied to plan across the full-year horizon.
-If the RO formulation is used, a decision-dependent uncertainty set is constructed from the historical samples (Section 5.3).
-High-consumption vehicles with infeasible trips are then filtered out (Section 5.4).
-Finally, chargers are planned through the proposed model in the deterministic or stochastic formulation (Section 5.5).
+Next, the parking window, effective parking duration, and energy consumption of each truck are extracted (Section 4.2).
+Then, hierarchical clustering based on the rearranged time-zone features is applied to the parking-duration samples to construct feature-based uncertainty sets (Section 4.3).
+High-consumption vehicles (HCVs) with infeasible trips are then filtered out (Section 4.4).
+Finally, the proposed model plans chargers using different scheduling operations (Section 4.5).
 4.1 Identifying potential EV charging points
-We used GPS records of 189 light-duty trucks at the mining site over one year.
+We used GPS data from 189 light-duty trucks at a mining site over one year.
 Their GPS data records include coordinates (i.e., longitude and latitude) and speed at each time step.
-We then grid-divided the entire mining site into 100-meter $\times$ 100 100-meter square regions and counted the stop frequency in each square.
+We then grid-divided the entire mining site into 100-meter $\times$ 100-meter square regions and counted the stop frequency in each square.
 We assumed that a truck stops if it moves less than 50 meters within a square region for at least five minutes.
-Comparing the number of stop frequencies in each square, we identified eight EV charging points with the highest number of stops throughout the year.
+By comparing the number of stop frequencies per square, we identified eight EV charging points with the highest annual stop frequency.
 Fig.
@@
 The detailed statistics of eight charging zones are listed in Table 10 in Appendix 8.1.
-The zone 1 (roadhouse) and zone 4 (equipment maintenance) have a large part of long duration time greater than 4 hours, which means that most of the overnight parking happens at zones 1 and 4.
+Zones 1 (roadhouse) and 4 (equipment maintenance) account for a large share of long-duration time greater than 4 hours, indicating that most overnight parking occurs in zones 1 and 4.
 Figure 5: Identified eight charging zones and roads connected to eight charging zones (indicated as gray dashed lines)
@@
 We expect, however, that this additional consumption could be offset with minor operational adjustments, such as modifying parking durations.
-5.3 Samples for constructing robust uncertainty sets
-To construct a robust uncertainty set for heterogeneous effective parking durations, we first collected all parking durations and parking window data across the year.
+4.3 ML clustering for constructing feature-based robust sets
+Figure 8: Parking duration distributions in a year in terms of four features: (a) truck category, (b) half-hour of a day, (c) zones, and (d) day of the month
+Figure 9: Average parking duration heatmap for the selected 20 trucks, constructed using rearranged zone–hour features.
+The heatmap displays only the first 10 hours for each of the eight zones, with zone boundaries indicated by black dashed lines.
+The hierarchical clustering dendrogram, based on feature distances, is shown on the left.
+To construct a feature-based, robust uncertainty set for heterogeneous parking duration, as illustrated in Section 2.7.1, we first collected all parking durations in a year.
 Fig.
-8 (a) illustrates the total time of actual parking duration versus the the total time of the full parking window per truck per day at eight zones.
-In each zone, a portion of the parking window cannot be used for charging, for example, when the truck is moving to find a parking spot.
+8 (a) - (d) illustrate parking duration distributions in terms of four features, truck category, half-hour of the day, zone, and day of the month.
+We observed correlations between truck categories, half-hour of the day, and zones and parking duration.
+We then rearranged the time-zone features based on their combinations, and this gives a total number of $8\text{zones}\times 48\text{time slots}=384$ features.
+We then apply the proposed hierarchy clustering to group trucks based on the re-arranged time-zone features, as demonstrated in Section 2.7.1.
+Before clustering, we normalized the parking duration of these trucks using z-scores.
+The z-score standardizes the data for each truck to a mean of zero with a standard deviation of one.
+In other words, the average parking duration of a truck is normalized to one, and the duration that is higher or lower than the average value is a positive or negative value, respectively.
 Fig.
-8 (b) presents the distribution of parking durations across eight zones, showing that zones 6, 7, and 8 exhibit greater variability compared to the others.
-Figure 8: (a) Actual parking duration versus full parking windows duration per day per truck and (b) box plots of actual parking duration distributions of eight zones
-Figure 9: (a) Mean and (b) variance of effective parking durations in minutes of 189 trucks by hour of the day in the charging zone 1
-We found that the actual parking duration at the mining site is strongly related to trucks, hour of the day, and charging zones, but has almost no correlation with day of the week or day of the month.
-This is because the mining site operates continuously with the same pattern throughout the year.
-We then calculate the mean and standard deviation of parking durations pertaining to trucks, hours of the day, and zones.
-Fig.
-9 shows the mean and variance of effective parking duration in minutes by hour of the day in the charging zone 1.
-Most trucks operating around the mining site have zero means and variances in their actual parking duration during the day.
+9 shows the normalized parking duration heatmap for the selected 20 trucks, constructed using rearranged zone–hour features, with the hierarchical clustering dendrogram on the left.
+We first grouped trucks by category.
+Using the category information, hierarchical clustering yields three subgroups per category, for a total of 24 subgroups.
+We also compare these methods with the other three: hierarchical clustering without category information, k-means, and grouping each truck separately.
+These comparisons and clustering results are detailed in Appendix 9.
 4.4 Set up
-We prepared three inputs based on the dataset of the year 2023, for the DS and IPS approaches, respectively.
-For the DS approach, we defined 10 statistical criteria, including total fleet distance, number of active trucks, and total fleet energy consumption.
+We prepared three inputs for the selective and full-year datasets for 2023.
+For the selective day approach, we defined 10 statistical criteria to select 39 days, including total fleet distance, number of active trucks, and total fleet energy consumption.
 Based on these criteria, we ranked all 365 days in a year and selected the top-ranking days.
-We experimented with different numbers of selected days, ranging from 10 to 40, and observed that charger installation results stabilized once the selection reached 30 days.
-Accordingly, we chose 39 days, including days adjacent to the top 10 ranked days, as some vehicles charge every two or three days, and representative sub-periods need to extend beyond battery cycles.
-The detailed selection criteria and procedures are provided in Appendix 9.2.
-For the IPS approach, we divided the full-year dataset into monthly subsets, incorporating overlapping days between consecutive months to maintain trip continuity.
-This procedure produced 12 monthly datasets, each spanning 34 days.
+We tested different numbers of selected days, from 10 to 40, and observed that charger installation results stabilized at 30 days.
+Accordingly, we selected 39 days, including days adjacent to the 10 top-ranked days, because some vehicles charge every 2 or 3 days, and representative sub-periods must extend beyond battery cycles.
+The detailed selection criteria and procedures are provided in Appendix 8.2.2.
+For the full year using rolling-horizon scheduling, we divided the full-year dataset into monthly subsets, with overlapping days between consecutive months to maintain trip continuity.
+This produces 12 monthly sub-datasets, each spanning 34 days.
 We conduct a preliminary simulation before optimization, in which each EV uses straightforward fast charging throughout the optimization horizon.
-In each simulation, a truck starts with a random SoC and is then fully charged at each parking site.
+In each simulation, a truck starts with a random SoC, and is then fully charged at each parking site.
 At the end of the optimization horizon, we track whether they could restore their original SoC.
-If the original SoC cannot be recovered, the truck is identified as a high consumption vehicle (HCV), which cannot be feasibly electrified if it sticks to the current operational scheduling.
-This also includes vehicles not passing the eight charging zones.
-The optimization is subsequently performed exclusively on the rest of the trucks, and we recorded the number of HCVs excluded from the optimization model in each simulation.
+If the original SoC cannot be recovered, the truck is identified as a HCV, which is not feasible to be electrified if they stick to the current operational scheduling.
+This also includes vehicles that do not pass eight charging zones.
+The optimization is subsequently performed exclusively on the remaining trucks, and we recorded the number of HCVs excluded from the optimization model for each simulation.
 All experiments were solved using the Gurobi commercial solver.
@@
 4.5 Result evaluation
+4.5.1 Performance metrics
+Apart from the number of fast and slow chargers installed, the following performance metrics are introduced to evaluate performance:
+  • Fast charging hours (FC Hrs, per truck per day): The aggregated duration, measured in half-hour intervals, that a truck uses a fast charger per day.
+  • Slow charging hours (SC Hrs, per truck per day): The aggregated duration, measured in half-hour intervals, that a truck uses a slow charger per day.
+  • Abandonment hours (Ab Hrs, per truck per day): The aggregated duration, measured in half-hour intervals, that a truck is abandoned and charged per day.
+  • Average charging power (Avg. CP, kW): The mean power charged during all charging time slots.
+  • Percentage of the number of periods when battery SoC is lower than 30% relative to the total time slots (Pct.
+of Low SoC, %): The total number of half-hour periods when the battery SoC is lower than 30%.
+  • Average battery SoC of each truck (Avg. SoC, %): The average battery SoC of each truck throughout the trip.
+  • Charging utilization rate (CUR, %): The average number of charging duration, measured in half-hour intervals, each charger experiences compared to the optimization horizon.
 4.5.2 The impacts of abandonment, range anxiety, and effective parking duration
-We first evaluate four deterministic cases to analyze the impacts of charging abandonment, range anxiety, and parking duration restrictions.
-Using the DS method, we draw on 39 selected representative days for the input datasets.
-The total energy consumption of the entire vehicle fleet for the 39-day optimization horizon is 83,050 kWh, which is around 11.33 kWh per day per truck.
-The total parking time of the entire vehicle fleet is 124,639 hours, which is around 17 hours per day per truck.
+Using 39 representative days, we evaluate the impacts of abandonment, range anxiety, and effective parking duration using the HS method.
   • Case 1 - Benchmark: Trucks are prohibited from waiting in any zone under normal circumstances.
-However, when a vehicle SoC drops to below 30%, it is permitted to pause for 30 minutes to recharge.
-Two designated zones (i.e., zone 1 and zone 4) accommodate the overnight charging with unlimited waiting time.
+However, when a vehicle’s SoC drops to below 30%, it is permitted to pause for 30 minutes to recharge.
+Two designated zones (i.e., zones 1 and 4) accommodate the overnight charging with unlimited waiting time.
 The model incorporates the specific parking duration for each parking window (i.e., $\frac{5}{30}\leq pp\leq 1$).
-  • Case 2 - Full parking time: This case maintains the same conditions as the benchmark case, except that it has no specific parking duration constraints for each parking window (i.e., $pp=1$).
+  • Case 2 - Full parking time: This case maintains the same as the benchmark case, except that it has no specific parking duration constraints for each parking window (i.e., $pp=1$).
   • Case 3 - No overnight charging: This case maintains the same conditions as the benchmark case, except that it has no special zones for the overnight charging.
-  • Case 4 - No range anxiety: This case maintains the same conditions as the benchmark case, except that it has no related constraints or penalty related to the range anxiety, in order words, there is no penalty for low SoC (i.e., $P^{low}=1$).
-Trucks are prohibited from waiting in any zone under normal circumstances, regardless of their SoC.
-In other words, the maximum waiting time is set to zero (i.e., $T^{max}_{i,d,t}=0$).
+  • Case 4 - No range anxiety: This case maintains the same conditions as the benchmark case, except that it has no related constraints or penalty related to the range anxiety, in order words, there is no penalty for low SoC (i.e., $P^{low}=0$).
+Trucks are prohibited from waiting in any zone under normal circumstances, regardless of their SoC; in other words, the maximum waiting time is set to zero (i.e., $T^{max}_{i,d,t}=0$).
+Figure 10: Zone-level differences in the numbers of slow and fast chargers for Cases 2–4 relative to the benchmark. Dashed bars represent fast chargers, while solid bars represent slow chargers.
 Table 3 summarizes optimization results in four deterministic cases under different assumptions about abandonment, range anxiety, and effective parking duration.
-The benchmark case (case 1) has a mixed type of chargers, which are 72 slow chargers and 7 fast chargers.
-Compared to the benchmark case, case 2 shows that a less number of fast chargers are needed if the model does not consider effective charging time in each parking time slot.
-In case 3, an additional number of 20 slow chargers are installed if there is no special zone allowing for overnight charging.
-In case 4, if the model does not consider range anxiety, seven additional slow chargers need to be installed to accommodate more frequent abandonment without waiting time.
-Instances
-# of
-HCVs
-# of slow /
-fast chargers
-Installation
-cost (\$)
-Scheduling
-penalty (\$)
-Gap (%) Time (h)
-Case 1 1 72/6 336,000 38,468 0.084 2.98
-Case 2 1 74/3 187,000 39,622 0.116 57.32
-Case 3 1 92/6 366,000 85,200 0.042 15.77
-Case 4 1 79/6 346,500 38,077 0.079 2.94
-Table 2: Summary of optimization results in four deterministic cases: benchmark, no parking time restriction, no overnight charging, and no range anxiety
-Figure 10: Zone-wise differences in the number of slow and fast chargers of cases 2-4 compared with the benchmark
+The benchmark case (case 1) has a mixed type of chargers, which are 72 of slow chargers and 6 of fast chargers.
 Fig.
-10 illustrates the differences in the numbers of fast and slow chargers in cases 2–4, compared with the benchmark.
-Across four cases, the number of installed chargers is proportional to the total duration of vehicle stops at each charging zone, as presented in Fig.
-8 (a), and zones 4 and 5 have the highest number of chargers.
-In case 1, setting the special zones does not relax the requirement of the fast charger installation in these zones, as in zones 1 and 4.
-In case 3, without considering the overnight charging, zones 1 and 4 will have a high number of slow chargers.
-In case 4, without considering the range anxiety, zones 5-7 will have a higher number of slow chargers, meaning vehicles passing through these three zones have a relatively low SoC than other zones.
-Instances
-Total slow
-charging
-hours
-Total fast
-charging
-hours
-Total
-abandonment
-hours
-Avg. charging
-hours
-(truck per day)
-Avg. charging
-power
-Case 1 3677.0 1352.5 26371.0 0.69 16.51
-Case 2 5779.0 555.5 26430.5 0.86 13.11
-Case 3 3493.5 1418.5 69353.0 0.67 16.90
-Case 4 3116.5 1248.5 26947.0 0.60 19.02
-Table 3: Summary of the scheduled charging hours, abandonment hours, and charging power across four deterministic cases
-Figure 11: The total number of (a) charging hours and (b) abandonment hours of the entire truck fleet in a day across four instances
-Figure 12: The total number of (a) charging hours and (b) abandonment hours of the truck fleet in different hours of a day in four deterministic cases
+10 illustrates differences in the numbers of fast and slow chargers in cases 2–4, compared to the benchmark.
+In the benchmark case, the number of installed chargers is proportional to the total duration of vehicle stops at each of the eight charging zones, as presented in Fig.
+8 (a), while zones 4 and 5 have the highest number of chargers among the eight zones.
+We then examine three cases in which overnight charging, range anxiety, and limited parking duration are not considered.
+Extending the parking duration to the full window reduces the number of fast chargers to three.
+Removing the designated overnight zones would require an additional 20 slow chargers to be installed in zones 1 and 4.
+In case 4, no range anxiety requires seven more slow chargers to be installed.
+Cases Names # of slow / fast chargers Installation costs (\$) Scheduling penalty (\$) Gap (%) Time (Hrs)
+1 Benchmark 72/6 336,000 38,468 0.084 2.98
+2 Full parking time 74/3 187,000 39,622 0.116 57.32
+3 No overnight charging 92/6 366,000 85,200 0.042 15.77
+4 No range anxiety 79/6 346,500 38,077 0.079 2.94
+Table 3: Summary of charging planning outcomes and overall scheduling penalties in four deterministic cases: Benchmark, Full parking time, No overnight charging, and No range anxiety
+Cases FC Hrs (truck $\cdot$ day) FCUR (%) SC Hrs (truck $\cdot$ day) SCUR (%) Ab Hrs (truck $\cdot$ day) Pct. low SoC (%) Avg. CP (kW)
+1 0.18 20.64 0.50 5.46 3.60 6.01 16.51
+2 0.08 19.78 0.79 8.34 3.60 4.91 13.11
+3 0.19 21.65 0.48 4.06 9.46 2.12 16.91
+4 0.17 19.06 0.43 4.21 3.68 62.57 19.03
+Table 4: Summary of all the performance metrics in four deterministic cases; The number in bold highlights the highest values among four cases
+Figure 11: The total number of (a) charging hours and (b) abandonment hours of the truck fleet in different zones in four deterministic cases
 Figure 12: (a) Average SoC of each truck and (b) the overall battery SoC for the entire fleet in four deterministic cases
-We calculate the charging power, charging hours, and abandonment hours across four deterministic cases, summarized in Table 3.
-The detailed scheduling solution of four cases are presented in Appendix 10.
-Case 2, which assumes full parking duration, results in the highest total charging hours but the lowest average charging power due to having the fewest fast chargers.
-Case 3, which excludes overnight charging, produces the greatest number of abandonment hours.
-Case 4, which removes range anxiety, leads to the fewest charging hours and the highest average charging power.
+Table 4 summarizes all the performance metrics in four deterministic cases, and the detailed scheduling solution of four cases are presented in Appendix 10.
+As expected, case 2, which assumes the full parking duration, has the highest fast-charging hours and utilization rate among the four cases.
+Case 3, which excludes overnight charging, has the highest number of fast-charging and abandonment hours.
+Case 4, with no range anxiety, leads to the highest percentage of low SoC events and the average charging power.
 Figs.
 11 (a) and (b) show the zone-wise comparison of charging and abandonment time, respectively.
-In case 2, with the full parking duration, zones 6 and 8 have a higher charging time.
+In case 2, with the full parking duration, zones 6 and 8 have longer charging times.
 In case 3, without the overnight charging, zones 1 and 4 have a significantly high number of abandonment hours.
-Figs.
-12 (a) and (b) show the total number of charging and abandonment hours of the total truck fleet in different hours of a day in four cases, respectively.
-Across four instances, the lowest number of charging hours appears at 12:00, and the highest number of charging hours appears at 22:00.
-If the model does not assign two special zones for overnight charging, the number of abandonment hours will increase by more than a factor of 2.
 Fig.
-13 (a) illustrates the average SoC of all trucks over 39 days, ranked from the highest to lowest, and Fig.
-13 (b) presents the overall battery SoC (maximum, minimum, and average) of the entire fleet in four cases.
+12 (a) illustrates the average SoC of all trucks over 39 days, ranked from the highest to the lowest, and Fig.
+12 (b) presents the overall battery SoC (i.e., the maximum, minimum, and average) of the entire fleet in four cases.
 Among four cases, case 4 (no range anxiety) yields the lowest average SoC ($28.4\%$), whereas case 2 with full parking duration achieves the highest average SoC ($45.3\%$).
-When the effective charging time constraint is applied, in other words, comparing case 2 to case 1 benchmark, part of the fleet experiences a sharp drop in the average SoC to very low levels ($\leq 20\%$), due to the limited charging duration in the parking windows.
-5.5.2 Robust infrastructure planning based on the whole year dataset
-[…]
-To ensure robust infrastructure planning throughout the year, we optimize the charging infrastructure using two solving approaches: the DS and IPS methods, as described in Section 3.
-For the DS method, we selected a 39-day horizon and imposed a maximum time limit of 40,000 seconds to solve each instance, recording the gap whenever the solver failed to converge within this limit.
-For the IPS method, we set a 6-hour time limit for each monthly instance.
-In addition, we applied the proposed fix-and-optimize heuristic to the 39-day instance under the DS framework, with the initial master problem aligned with the deterministic benchmark case.
-Instances
-($\sigma$, $\gamma$)
-Solving
-methods
-# of
-HCVs
-# of slow/fast
-chargers
-Installation
-costs (\$)
-Scheduling
-penalty (\$)
-Gap (%)
-Time (h)
-(0, 1) DS 19 68/6 330,000 35,678 9.06 11.1
-DS - h 19 68/6 330,000 35,679 0.07 2.40
-IPS 48 77/5 305,500 33,466 0.09 11.3
-(1, 1) DS 19 68/6 330,000 35,658 9.04 11.1
-DS - h 19 69/6 331,500 36,065 0.05 2.15
-IPS 48 75/7 378,500 33,331 0.05 12.0
-(2, 1) DS 19 68/6 330,000 35,736 9.02 11.1
-DS - h 19 69/6 331,500 35,863 0.05 5.20
-IPS 48 76/8 418,000 33,402 0.05 23.8
-(2, 0.5) DS 19 69/5 293,500 35,931 0.06 7.42
-DS - h 19 68/6 330,000 35,715 0.06 3.29
-IPS 48 74/8 415,000 22,541 0.03 29.8
-Table 4: Summary of optimization results across four instances using DS, DS with heuristic (DS-h), and IPS methods.
-For the IPS method, the reported gap is the optimal gap of the final instance in the iterative planning
-Figure 14: The total number of chargers in each month using the IPS method in four instances (in colored lines) versus the total number of trucks parking and charged (in gray bars) in all eight zones
-[…]
-Figure 15: Zone-wise differences in the number of slow and fast chargers in four stochastic cases compared to the benchmark
-When comparing the three solution methods, the IPS method yields the highest number of chargers across all four instances using the full-year datasets.
-The DS and DS-h methods have an even lower number of chargers compared with the deterministic benchmark case, since an additional 18 trucks are excluded as HCVs due to the short parking duration.
+When the effective charging time constraint is applied, in other words, comparing the benchmark to case 2, part of the fleet in the benchmark case experiences a sharp drop in the average SoC to very low levels ($\leq 20\%$), due to the limited charging duration in the parking windows.
+4.5.3 The impacts of different RO-DDU formulations for parking duration uncertainty
+We next evaluate the performance of two RO-DDU formulations for the charging power constraint under parking duration uncertainty as in Section 2.7.2.
+These formulations differ in whether the uncertainty set depends affinely on instantaneous charging decisions (hereafter referred to as ‘Instantaneous’) or on cluster-level expected charging decisions (hereafter referred to as ‘Expect’).
+We adopt fleet groups obtained from hierarchical clustering that incorporates truck category information, as described in Section 4.3.
+We then apply different scaling factors for slow chargers ($k_{\text{slow}}=4$) and fast chargers ($k_{\text{fast}}=2$) in the instantaneous and cluster-level RO-DDU formulations.
+Here, we envisage a pessimistic scenario in which parking duration will be reduced after BEV chargers are installed.
+In addition, we include an RO independent of charging decisions (‘RO-DI’) as a benchmark.
+All three cases are evaluated using 39 selected days, as defined by the IPS method.
+After obtaining the charger planning results, we use the sample average approximation (SAA) to compare the scheduling performance of the Instantaneous and Expect cases.
+Cases Formulations # of HCVs # of slow / fast chargers Installation costs (\$) Scheduling penalty (\$) Gap (%) Time (Hrs)
+1 RO - DI 19 67 / 6 328,500 20,911 0.01 0.48
+2 Instant. 19 67 / 6 328,500 30,400 0.01 0.52
+3 Expect 19 70 / 6 333,000 31,306 0.01 10.21
+Table 5: Summary of charging planning outcomes and scheduling penalties under three RO formulations (low-penalty case)
+Figure 13: Zone-wise differences in the numbers of slow and fast chargers for three RO cases: RO - DI, RO-DDU (Instantaneous), and RO-DDU (Expect). The benchmark case is the RO - DI case
+Sample Cases FC Hrs (truck $\cdot$ day) SC Hrs (truck $\cdot$ day) Ab Hrs (truck $\cdot$ day) Pct. of low SoC (%) Avg. CP (kW) Tot. Penalty (\$/truck $\cdot$ day)
+In sample RO-DI 0.16 0.45 4.14 1.52 9.47 4.43
+Instant. 0.22 0.52 4.13 1.72 7.80 4.51
+Expect. 0.16 0.46 4.10 2.35 9.48 4.40
+Out of sample Instant. 0.21 0.39 3.77 4.38 9.28 4.02
+Expect. 0.21 0.39 3.77 4.90 9.33 4.01
+Table 6: Summary of performance metrics for in-sample and out-of-sample evaluations under RO-DI, instantaneous DDU, and expectation-based DDU at the low-penalty case.
+Values are rounded to two decimal places.
+The values in bold indicate the highest value within each sample group and column.
+Table 5 summarizes the charger installation results of three RO cases obtained in the in-sample datasets.
+Using the IPS approach can decompose the problem and solve three instances quickly, compared to the HS method.
+It relaxes the battery energy balance over a long scheduling horizon.
 Fig.
-14 shows the total number of chargers (i.e., the sum of fast and slow chargers) determined by the IPS procedure from monthly subproblems over 12 months in the four instances.
-From January to December, the number of trucks parking and charged in any of the eight zones rises from 156 to 182, leading to a steady increase in the required charger installations.
-As the parking duration uncertainty grows, the total number of EV chargers—particularly fast chargers—also increases.
-In the fourth instance, a larger share of fast chargers is installed compared to the third instance, since fast chargers are less affected by uncertainty than slow chargers.
-Across four stochastic instances, the only difference between the DS and DS-h results lies in the installation or placement of a single fast charger.
-This indicates that the assignment of specific fast chargers can hinder the convergence of the DS method.
-In contrast, the proposed heuristic prioritizes placing the fast charger, allowing the model to achieve the optimal solution more efficiently.
-The detailed solution process using the DS-h method is presented in Appendix 11.
+13 shows their zone-wise differences in the numbers of installed slow and fast chargers.
+Only when using the cluster formulation, the required number of chargers increases, particularly in zones 6, 7, and 8, where the charging duration variability is higher than in other zones.
+In addition, the cluster formulation requires substantially longer solution times due to the introduction of nonlinear charging-power constraints.
+Table 6 summarizes the scheduling performance metrics of in-sample and out-of-sample tests in three cases.
+During in-sample tests, the RO-DI formulation yields the highest charging power, followed by RO-DDU (Expect) and RO-DDU (Instantaneous).
+This aligns with the theoretical analysis of charging power bounds via cluster-level aggregation in Section 7.3.
+The RO-DDU (expect) formulation accounts for decision dependence and the correlation among charging decisions within the same cluster.
+Therefore, the RO-DDU (expect) formulation has the highest scheduling penalty among the three cases.
+In out-of-sample tests, the scheduling penalty obtained from the RO-DDU (expect) formulation becomes less than the RO-DDU (Instantaneous) formulation.
+This means that, by accounting for correlations within the cluster, the expected formulation yields more efficient scheduling.
+Data Range $\Delta_{c}$ Solving methods # of HCVs # of slow/fast chargers Installation costs (\$) Scheduling penalty (\$) Gap (%) Time (Hrs)
+DS 0 HS 19 68/6 330,000 35,678 9.06 11.1
+HS - h 19 68/6 330,000 35,679 0.07 2.40
+DS $\pm\sigma$ HS 19 68/6 330,000 35,658 9.04 11.1
+HS - h 19 69/6 331,500 36,065 0.05 2.15
+DS $\pm\sigma$ HS 19 67/8 442,500 35,565 23.19 11.1
+DDU HS - h 19 70/6 330,000 22,455 0.06 3.29
+DS $\pm 2\sigma$ HS 19 68/6 330,000 35,736 9.02 11.1
+HS - h 19 69/6 331,500 35,863 0.05 5.20
+Table 7: Summary of optimization results with the varying size of uncertainty set based on the selected 39-day and full-year dataset using HS, HS with heuristic (HS - h)
+Tackling high computation costs with the ‘fix-and-optimize’ heuristic: Considering the high computation costs of the 39-day instance under HS scheduling, we leverage the fix-and-optimize heuristic to solve problems.
+We impose a maximum time limit of 40,000 seconds for each instance and report the optimality gap when the solver fails to converge within this limit.
+In addition, we initialize the master problem as the warm start.
+As presented in Table 7, we found that solving these HS instances directly is difficult, with all four cases reaching the maximum time limit.
+Using the proposed heuristic method (HS - h), the problem can reach an acceptable optimality gap of less than 0.1% while keeping the computation time within the prescribed limits.
+We used a warm start because the solution to the deterministic benchmark case is available, and completing 10 iterations takes approximately one hour.
+Across cases with varying uncertainty sets, the only difference between the HS and HS-h methods is the installation or placement of a single fast charger.
+This indicates that assigning specific fast chargers can hinder convergence.
+In contrast, the proposed heuristic prioritizes placing the fast charger, allowing the model to converge more efficiently.
+The detailed solution processes using the HS-h method are presented in Appendix 11.
+4.5.4 Robust infrastructure planning based on the full year dataset
+To ensure robust infrastructure planning over the year, we use the full-year dataset to plan the charging infrastructure under the IPS methods, as described in Section 3.1.
+These three full-year instances, corresponding to the size of uncertainty set increasing from 0 to $2\sigma$, provide the most robust planning results compared with the selective datasets.
+Data Range $\Delta_{c}$ Solving methods # of HCVs # of slow/fast chargers Installation costs (\$) Scheduling penalty (\$) Gap (%) Time (Hrs)
+FY 0 IPS 48 77/5 305,500 33,466 0.09 11.3
+$\pm\sigma$ IPS 48 75/7 378,500 33,331 0.05 12.0
+$\pm 2\sigma$ IPS 48 76/8 418,000 33,402 0.05 23.8
+Table 8: Summary of optimization results with the full-year dataset using the IPS method. For the IPS method, the reported gap is the optimality gap of the final period in the iterative planning.
+Figure 14: The total number of chargers in each month based on the full-year dataset using the IPS method in three instances with varying sizes of uncertainty sets (in colored lines) versus the total number of trucks parking and charged (in gray bars) in all eight zones
+Table 7 summarizes the optimization results.
+When extending the data range to the full year using the IPS method, between 1 and 3 iterations exceeded the time limit.
+However, all three full-year instances converge by the end of the 12-month horizon.
+In the final instance involving joint planning and scheduling, an optimality gap of less than 0.1% is achieved.
+Fig.15 shows zone-wise differences in the number of slow and fast chargers in three full-year cases with the increasing size of uncertainty sets.
+As uncertainty increases, two or three additional fast chargers are replaced with slow chargers.
+Figure 15: Zone-level variations in the number of slow and fast chargers across three full-year cases are illustrated.
+The benchmark corresponds to the case with a zero-sized uncertainty set, in which parking duration is fixed at the mean value for each cluster.
+When comparing results from the 39-day dataset with those from the full-year dataset, the latter yields more chargers because the full-year data captures the end-of-year period with the highest vehicle counts.
+In contrast, the results derived from the 39-day dataset require even fewer chargers than the deterministic benchmark, because an additional 18 trucks are excluded from the HCV set due to their very short parking durations.
 Fig.
-15 shows the zone-wise differences in the number of slow and fast chargers when comparing four stochastic instances to the benchmark case (i.e., case 1 in the deterministic formulation).
-In general, zones 1, 2, 4, and 7 have an increase in the number of slow chargers, while zone 6 has a reduction.
-It is because that zone 6 has a quite high variance in the parking duration.
-The model optimizes more chargers and schedules more trucks to be charged at zones 1 or 2, which are connected to zone 6, in four stochastic cases.
-5.5.3 The impacts of uncertain parking duration
-To evaluate the impact of uncertain parking duration on the charger planning and scheduling, we first simulated the monthly schedule based on the charger planning results for the case where $\sigma=2$ and $\gamma=0$, using the DS-h and IPS methods, respectively.
-We found that the charging installation obtained from the DS-h method meets the requirements for only six months, where the number of trucks dispatched is less, whereas those obtained from the IPS method satisfy the requirements for all 12 months.
-This discrepancy arises from the steadily increasing trend of truck parking and charging activities throughout the year, which the day-selection approach fails to capture adequately during peak months.
-Instances
-($\sigma$, $\gamma$)
-Avg. slow
-charging
-hours
-Avg. fast
-charging
-hours
-Avg.
-abandonment
-hours
-Avg. charging
-hours
-(truck per day)
-Avg. charging
-power (kW)
-(0, 1) 3757.6 1280.7 20212.9 0.83 8.78
-(1, 1) 2166.9 1249.9 20229.9 0.56 12.95
-(2, 1) 2126.3 1357.5 20155.1 0.57 12.70
-(2, 0.5) 515.1 1033.0 20191.6 0.25 28.59
-Table 5: Summary of the average scheduled charging hours, abandonment hours, and charging power in a year across four instances
-Figure 16: Boxplots of the monthly slow and fast charging hours across four stochastic instances versus the average of charging hours by month
-We then evaluated four stochastic instances using the IPS method by simulating the charging scheduling in 12 months and calculated the average value of the total charging and abandonment hours by month.
+14 shows the total number of chargers (i.e., the sum of the number of fast and slow chargers) determined by the IPS procedure from monthly subproblems over 12 months in three instances.
+From January to December, the number of trucks parked and charged across the eight zones increases from 156 to 182, resulting in a steady rise in the number of required charger installations.
+As the parking duration uncertainty grows, the total number of EV chargers increases, particularly for fast chargers.
+We then evaluated three full-year instances by calculating their performance metrics across 12 months.
+Table 9 summarizes the performance metrics of all three cases.
+As the uncertainty increases, both fast and low charger utilization rates reduce.
+The average abandonment hour remains constant across the three stochastic instances.
+$\Delta_{c}$ FC Hrs (truck $\cdot$ day) FCUR (%) SC Hrs (truck $\cdot$ day) SCUR (%) Ab Hrs (truck $\cdot$ day) Pct. low SoC (%) Avg. CP (kW)
+0 0.21 31.39 0.62 5.98 6.64 6.97 12.44
+$\pm\sigma$ 0.21 21.88 0.36 3.54 6.64 2.46 13.18
+$\pm 2\sigma$ 0.22 20.80 0.35 3.43 6.64 3.59 12.80
+Table 9: Summary of all performance metrics in three stochastic cases. The number in bold highlights the highest values among the three cases.
+Figure 16: (a) Distributions of monthly slow and fast charging hours of the entire truck fleet and (b) the monthly charging hour trends across a year in three uncertainty scenarios
 Fig.
-16 shows the box-plots of the total number of slow and fast charging hours by months in four stochastic instances using the IPS method, and Table 5 summarizes the average scheduled charging hours, abandonment hours, and charging power by months.
-As the uncertainty increases where the standard deviation of the parking duration $\sigma$ increases from 0 to 2, the fast charger usage rises while the slow charger usage declines, since the shorter parking durations incentivize drivers to rely more on the fast charging.
-As the fast chargers are more frequently used, the average charging power increases from 8.78 kW when the standard deviation equals 0, to 12.95 kW and 12.70 kW when the standard deviation equals 1 and 2, respectively.
-The average abandonment hour keeps the same level across four stochastic instances.
-For the second and third instances, where the standard deviations are 1 and 2, respectively, the third instance exhibits longer total charging hours and lower charging rates, even though fast-charging hours are greater than in the second instance.
-This is because the increased uncertainty in parking duration prolongs overall charging time, offsetting the benefit of more frequent fast charging.
-In contrast, in the fourth instance—where fast chargers effectively mitigate parking duration uncertainty—the total charging time decreases significantly, with trucks relying on fast chargers for most of the period.
-The average monthly slow-charging time is only 515.1 hours.
+16 (a) shows distributions of monthly slow and fast charging hours of the entire truck fleet in the three stochastic instances, and Fig.
+16 (b) shows the total monthly charging hour trends.
+By comparing three cases, we have the following observations: (1) With the increasing uncertainty of the charging duration, the total charging hours (the parking slots being decided to charge) reduce, but the fast charging hours increase.
+This indicates that the scheduling system seeks to avoid parking slots with highly uncertain parking durations and instead uses fast chargers more frequently.
+(2) If we further increase the uncertainty of charging duration, the total charging hours will increase, which is opposite to the previous trend, as more charging slots are required to meet the energy consumption.
 4.6 Conclusion and discussion
-We developed a two-stage robust charging infrastructure planning model for the electrification of light-duty trucks at industrial sites with mixed-type chargers with different charging power.
-The first stage plans the charging infrastructure based on the selected charging locations and charging types.
-The second stage optimizes the charging schedules considering different waiting times, charging abandonment, and range anxiety.
-A case study is conducted on an open-pit mining site with eight potential charging zones.
-We used three input datasets extracted from the whole-year GPS data from around 200 trucks: (i) parking indications at each half-hour interval, (ii) actual parking durations of each parking window, and (iii) half-hour energy consumption of all trucks.
-5.6.1 The impacts of overnight charging and range anxiety on charging behaviors
-Among eight zones, we assume that two designated zones allow overnight parking; otherwise, trucks will forgo charging if no chargers are available, unless their battery level falls below 30%.
+We develop a robust model for planning and scheduling the charging infrastructure required to electrify light-duty trucks at industrial sites with mixed charger types.
+The model determines charging infrastructure deployment by selecting charging locations and charger types, and jointly optimizes charging schedules while accounting for waiting time, charging abandonment, and range anxiety.
+We conduct a case study at an open-pit mining site comprising eight charging zones.
+Three input datasets are constructed from year-long GPS records of approximately 200 operational trucks: (i) parking time slots discretized into half-hour intervals, (ii) realized parking durations within each half-hour window, and (iii) half-hourly energy consumption of all trucks.
+We assume fixed charging efficiencies for both slow and fast chargers and do not account for energy consumption during parking periods.
+4.6.1 No overnight charging or no range anxiety will significantly increase the charger installations
+Among eight zones, we assume that two designated zones allow overnight parking; otherwise, trucks will forgo charging if no chargers are available.
 The battery SoC status below 30% is assumed to trigger the range anxiety, resulting in an additional half-hour waiting time.
-The benchmark case based on the selected 39 days yields 72 slow chargers and 6 fast chargers.
-We then examine three cases where overnight charging, range anxiety, and limited parking duration in each parking window are not considered, respectively.
-Extending the parking duration to the full window reduces the number of fast chargers to 3.
-Removing the designated overnight zones leads to a dramatic increase in abandonment hours by 1.63 times and an additional 20 slow chargers installed in total.
-No range anxiety requires seven more slow chargers to be installed.
-In addition, we found that the charging hours only occupy around 4% of the total parking time, meaning the limited waiting time and consequential abandonment are the main reasons that lead to a number of slow chargers at industrial sites.
-5.6.2 Robust charging infrastructure planning considering uncertain parking duration
-To utilize the full-year dataset, we proposed two methods—the DS and IPS approaches—for charging infrastructure planning.
-The uncertain parking duration within each half-hour parking window is represented using a robust uncertainty set, whose bounds depend on the charging decisions.
-The IPS method ensures robust scheduling performance throughout the year by planning for a larger number of chargers, whereas the DS method fails in several months when the charging demand peaks.
-When the uncertainty of charging duration increases (i.e., the lower bound of the uncertainty set decreases), the total number of fast-charging hours rises, while the total number of slow-charging hours declines.
-Consequently, the overall average charging power increases under the mixed-type charging infrastructure.
-However, further reducing the lower bound of the robust set for parking duration to the minimum level (i.e., 5 minutes) leads to longer total charging hours and a decrease in average charging power.
-If the model incorporates the mitigating effect of fast charging on parking duration uncertainty—observed in many real-world studies—a smaller number of slow chargers would be planned and utilized.
-5.6.3 Heuristic methods to assist the robust scheduling with a long optimization horizon
-The proposed two-stage planning and scheduling framework is formulated as a large-scale MILP model.
-For a 39-day scheduling instance with 189 trucks, the model contains over 1.5 million binary variables without pre-solving, making it computationally challenging.
-To address this, we introduce a fix-and-optimize heuristic for the charging scheduling under uncertainty.
-This is to forecast the high charging demand due to reduced parking duration and to pre-assign fast charging actions in the scheduling.
+We found that neither overnight charging nor range anxiety significantly increases the number of charger installations, as both behaviors entail additional waiting time and reduce the need for installations.
+In addition, the charging hours account for only about 4% of the total parking time.
+This means limited waiting time and resulting abandonment are the main reasons for the number of charger installations at industrial sites.
+4.6.2 A greater uncertainty in charging duration leads to changes in total charging hours as well as shifts in charging type
+We assume that parking duration in the parking window is stochastic.
+To identify patterns, hierarchical clustering is used to group trucks.
+Then, the uncertain parking duration of each truck can be represented as a robust, feature-based uncertainty set.
+As charging duration uncertainty increases, total charging hours decline, reducing utilization rates for both slow and fast chargers.
+Increased uncertainty also shifts charging patterns, with total fast-charging hours increasing while slow-charging hours decrease.
+However, when the lower bound of the robust parking-duration set is further reduced to its minimum value (i.e., 5 minutes), the total charging hours increase to meet each truck’s energy requirements.
+4.6.3 Alternative decision-dependent formulations that rely on clustering produce different charger planning results
+We consider that the uncertainty of charging duration is decision-dependent, which impacts and is impacted by charging decisions (i.e., whether or not charging and charging types).
+We then investigate two decision-dependent formulations, assuming an affine dependence on instantaneous and cluster-level expected decisions.
+We found that cluster-level expected decisions yield tighter bounds on scheduling and a higher charging power level when we assume that charging decisions reduce parking duration.
+The instantaneous formulation always imposes a certain risk on individual parking time slots and neglects aggregated, correlated charging/parking behaviors across the fleet.
+4.6.4 Fix-and-optimize heuristic can assist the robust scheduling depending on the forecasts
+The proposed planning and scheduling framework is formulated as a large-scale MILP model.
+For a 39-day instance with 188 trucks, the model contains over 1.5 million binary variables during pre-solving, making it computationally challenging.
+To address this, we introduce a fix-and-optimize heuristic for scheduling the charging operations under uncertainty.
+This is to forecast high charging demand due to reduced parking durations and to schedule fast-charging actions.
 This reduces the computation time by approximately 2-5 times for 39-day instances and yields the optimality gap of the problem below 0.1%.
@@
 For example, future work could incorporate additional charging options, such as multiple types of fast chargers with rated powers exceeding 150 kW, to further reduce the number of required HCVs.
-Moreover, charging behavior modeling could be improved by integrating empirical data from existing electrified fleets to account for factors like variations in range anxiety at different times of day.
+Moreover, charging behavior modeling could be improved by integrating empirical data from existing electrified fleets to account for factors such as variations in range anxiety across times of day.
 4.7 Acknowledgment